### v1

# Characterization and Decidability of FC-Definable Regular Languages

Abstract
$\mathsf{FC}$ is a first-order logic that reasons over all factors of a finite word using concatenation, and can define non-regular languages like that of all squares ($ww$). In this paper, we establish that there are regular languages that are not $\mathsf{FC}$-definable. Moreover, we give a decidable characterization of the $\mathsf{FC}$-definable regular languages in terms of algebra, automata, and regular expressions. The latter of which is natural and concise: Star-free generalized regular expressions extended with the Kleene star of terminal words.
Index Terms:
Finite model theory, first-order logic, regular languages.
I Introduction
The logic $\mathsf{FC}$ was introduced by Freydenberger and Peterfreund [1] as a new approach to first-order logic over finite words. Commonly, logic treats words as a sequence of positions which are given symbols with symbol predicates (e.g., see [2]). Instead, $\mathsf{FC}$ reasons over the set of factors of an input word using a concatenation relation $x\mathbin{\dot{=}}y\cdot z$ along with constant symbols for terminal symbols and for the empty word. For example, $\exists x\exists y\exists z\colon(x\mathbin{\dot{=}}y\cdot y)\land(y\mathbin{\dot{=}}\mathtt{b}\cdot z)$, with $\mathtt{b}$ being a terminal symbol, defines the language of those words that contain a factor of the form $\mathtt{b}v\mathtt{b}v$ with $v\in\Sigma^{*}$. $\mathsf{FC}$ is a finite model variant of the theory of concatenation (introduced by Quine [3]). Both combine word equations (which have been extensively studied [4, 5, 6, 7, 8]) with first-order logic.
A motivation for $\mathsf{FC}$ is the connection to document spanners, which were introduced by Fagin, Kimelfeld, Reiss and Vansummeren [9]. Document spanners query text documents by first extracting tables using regular expressions, and then applying a relational algebra to those extractions.
$\mathsf{FC}$ extended with regular constraints (atomic formulas which ensure that a variable is replaced with a word from a certain regular language) has the same expressive power as the generalized core spanners [1], a particular class of document spanners. Moreover, the existential-positive fragment of $\mathsf{FC}$ with regular constraints [^1] Although, the existential-positive fragment requires an extra constant symbol, which Freydenberger and Peterfreund [1] call the universe variable. has the same expressive power as the core spanners [1] – which were introduced by Fagin et al. [9] to capture the core functionality of IBM’s Annotation Query Language. This strong connection between $\mathsf{FC}$ and document spanners allows one to bring techniques from finite model theory to text querying, such as formulas with bounded width [1], acyclic conjunctive queries [10], and Ehrenfeucht-Fraïssé games [11]. However, the use of regular constraints naturally raises the question as to whether they are needed.
This paper shows that the class of $\mathsf{FC}$-definable regular languages is a proper subset of the regular languages (over non-unary alphabets [^2] The situation is more straightforward for unary alphabets: $\mathsf{FC}$ over a unary alphabet is exactly the semi-linear languages [11] and thus can define all the unary regular languages [12].). This demonstrates that an extension by regular constraints is necessary when using $\mathsf{FC}$ as a logic for document spanners. Moreover, stepping aside from the database theory aspects of $\mathsf{FC}$, this paper provides a comprehensive answer to a fundamental question regarding the expressive power of first-order logic with concatenation. It follows in the long-standing tradition of characterizing the regular languages definable in various logics (see Straubing [2] as a starting point), focusing on the logic $\mathsf{FC}$. The main results provide a decidable characterization of the $\mathsf{FC}$-definable regular languages in terms of (generalized) regular expressions, automata, and algebra.
Regular Expression Characterization
Arguably, the most natural and concise formulation of the $\mathsf{FC}$-definable regular languages is the regular expression formulation: Take the star-free generalized regular expressions (which are usual regular expressions without Kleene star, but with complement) and add the Kleene star of terminal words. For example, $\bigl{(}\mathtt{abab}\cdot(\mathtt{abab})^{*}\bigr{)}^{c}$, where the superscript $c$ denotes complement, describes the language of words that are not of the form $(\mathtt{ab})^{2n}$ for any $n>0$. To formalize this class, we use the star-free closure operator; introduced by Place and Zeitoun in [13]. However, we rely on the results and formulations given in the more recent version [14]. The star-free closure takes a class of regular languages $\mathcal{C}$, and maps it to $\mathsf{SF}(\mathcal{C})$ which is the smallest class containing $\mathcal{C}$, the singletons $\{\mathtt{a}\}$ for terminal symbols, and is closed under concatenation, union, and complement. A main result of this paper shows that the class of $\mathsf{FC}$-definable regular languages is exactly $\mathsf{SF}(\mathcal{R})$ where :=$\mathcal{R}:=\{w^{*}\mid w\in\Sigma^{*}\}$. Using the notion of star-free closure allows us to draw upon the results and techniques of [13, 14], in particular, the algebraic characterization of the star-free closure of so-called prevarieties.
Automata Characterization
The automata characterization is given by a necessary and sufficient criterion relying on a new notion which we call a loop-step cycle. A minimal DFA has a loop-step cycle if there are two words $w$ and $v$, that are not repetitions of the same word, and a sequence of $n\geqslant 2$ pairwise distinct states $p_{0},p_{1},\dots,p_{n-1}$ such that reading $w$ in any state $p_{i}$ results in again being in state $p_{i}$, and reading $v$ in any state $p_{i}$ result in moving to state $p_{i+1\pmod{n}}$. Notice that the minimal DFA for the language of words with an even number of $\mathtt{a}$ symbols (over the alphabet $\{\mathtt{a},\mathtt{b}\}$) has a loop-step cycle (see Fig. 1). A main result of this paper shows that a regular language is $\mathsf{FC}$-definable if, and only if, its minimal DFA does not have a loop-step cycle. We shall use this automata characterization to show that it is decidable (and $\mathsf{PSPACE}$-complete) to determine whether a regular language (given as a minimal DFA) is $\mathsf{FC}$-definable.
$p_{0}$$p_{1}$$\mathtt{a}$$\mathtt{a}$$\mathtt{b}$$\mathtt{b}$
Figure 1: The minimal DFA for the language of words with an even number of $\mathtt{a}$ symbols.
Algebraic Characterization
Every regular language $L$ is associated with a finite monoid $M_{L}$ known as the syntactic monoid [^3] For this introduction, the precise definition of $M_{L}$ is not important., and a function $\eta_{L}$ called the syntactic morphism which maps words to elements of this monoid (i.e., $\eta_{L}\colon\Sigma^{*}\rightarrow M_{L}$), see Pin [15] for example. Then, we have that $L=\eta_{L}^{-1}(F)$ for some $F\subseteq M_{L}$. A main result of this paper characterizes the $\mathsf{FC}$-definable regular languages by a class of syntactic morphisms which we call group primitive morphisms. A syntactic morphism $\eta_{L}\colon\Sigma^{*}\rightarrow M_{L}$ is group primitive if the inverse of any periodic element $m\in M_{L}$ – that is, where $m^{n}\neq m^{n+1}$ for all $n\in\mathbb{N}_{+}$ – is a subset of $w^{*}$ for some $w\in\Sigma^{*}$.
Related Work
We refer to the more common first-order logic over strings (a linear order with symbol predicates) by $\mathsf{FO}[<]$. Our characterization of the $\mathsf{FC}$-definable regular languages parallels the characterization of $\mathsf{FO}[<]$-definable regular languages [16, 17]. Each of the formalisms of our characterization naturally generalizes the characterizations of the $\mathsf{FO}[<]$-languages: The class $\mathsf{SF}(\mathcal{R})$ extends the star-free languages [17]. The notion of automata with loop-step cycles is analogous to finite-automaton cycle existence [18]. Lastly, group primitive languages generalize the languages definable by aperiodic monoids [17].
Regarding the expressive power of $\mathsf{FC}$, Freydenberger and Peterfreund [1] showed that $\{\mathtt{a}^{n}\mathtt{b}^{n}\mid n\geqslant 0\}$ is not $\mathsf{FC}$-definable. Then, using Ehrenfeucht-Fraïssé games, Thompson and Freydenberger [11] provided a general tool for $\mathsf{FC}$ inexpressibility called the fooling lemma. Conjunctive query fragments of $\mathsf{FC}$ were considered by Thompson and Freydenberger [19], where the expressive power was compared to other language generators.
Structure of Paper
Section II gives some preliminary definitions. We start Section III by giving the formal definitions of star-free closure (Section III-A), group primitive languages (Section III-B), and loop-step cycles (Section III-C); all of which are required for formulating our main result: a decidable characterization of the $\mathsf{FC}$-definable regular languages (Section III-D). The rest of the article is dedicated to proving this characterization. At the end of Section III-D, we give an overview of the proof structure of the main result. We close the paper in Section VIII. Some technical details have been deferred to the appendix.
II Preliminaries
Let :=$\mathbb{N}:=\{0,1,2,\dots\}$ and let +:=$\mathbb{N}_{+}:=\mathbb{N}\setminus\{0\}$ where $\setminus$ denotes set difference. The cardinality of a set $S$ is denoted by $|S|$. For $n\in\mathbb{N}_{+}$, we use $[n]$ for $\{i\in\mathbb{N}\mid 1\leqslant i\leqslant n\}$. For a vector $\vec{a}\in A^{k}$, for some set $A$ and $k\in\mathbb{N}$, we write $x\in\vec{a}$ to denote that $x$ is a component of $\vec{a}$. For a finite set $A\subset\mathbb{N}$, the minimum and maximum elements of $A$ are denoted by $\mathsf{min}(A)$ and $\mathsf{max}(A)$ respectively.
Words and Languages
We use $\Sigma$ for a fixed and finite alphabet of terminal symbols of size $|\Sigma|\geqslant 2$. We write $\Sigma^{*}$ for the set of all words of finite length built from symbols in $\Sigma$, we let $\varepsilon$ denote the empty word, and we let $\Sigma^{+}:=\Sigma^{*}\setminus\{\varepsilon\}$. For a word, $w\in\Sigma^{*}$ we write $w^{*}$ for the language $\{w^{n}\mid n\in\mathbb{N}\}$, where $w^{n}$ is $n$ consecutive repetitions of $w$.
If $w=w_{1}\cdot w_{2}\cdot w_{3}$ where $w,w_{1},w_{2},w_{3}\in\Sigma^{*}$, then $w_{1}$ is a prefix of $w$ (denoted $w_{1}\sqsubseteq_{\mathsf{pref}}w$), $w_{2}$ is a factor of $w$ (denoted $w_{2}\sqsubseteq w$), and $w_{3}$ is a suffix of $w$ (denoted $w_{3}\sqsubseteq_{\mathsf{suff}}w$). If $w_{2}\neq w$ also holds, then $w_{2}\sqsubset w$ and we use the analogous symbols $\sqsubset_{\mathsf{suff}}$ and $\sqsubset_{\mathsf{pref}}$. If $w_{1}\neq\varepsilon\neq w_{3}$, then we call $w_{2}$ an internal factor of $w$. The set of all factors of $w\in\Sigma^{*}$ is denoted by $\mathsf{facts}(w):=\{u\in\Sigma^{*}\mid u\sqsubseteq w\}$. We use $|w|$ for the length of $w\in\Sigma^{*}$; and for some $\mathtt{a}\in\Sigma$, we use $|w|_{\mathtt{a}}$ to denote the number of occurrences of $\mathtt{a}$ within $w$.
A word $w\in\Sigma^{+}$ is primitive if for any $u\in\Sigma^{+}$ and $n\in\mathbb{N}$, we have that $w=u^{n}$ implies $n=1$. In other words, $w$ is not a repetition of a smaller word. A word $w\in\Sigma^{+}$ is imprimitive if it is not primitive. For any word $w\in\Sigma^{+}$, the primitive root of $w$ is the unique primitive word $\varrho(w)\in\Sigma^{+}$ such that $w=\varrho(w)^{k}$ for some $k\geqslant 1$. For a language $L\subseteq\Sigma^{*}$, let $\varrho(L):=\{\varrho(w)\mid w\in L\setminus\{\varepsilon\}\}$.
Algebraic Concepts
A monoid $(M,\cdot,e)$ is a set $M$ with an associative multiplication operation $\cdot$ and an identity element $e$. When the multiplication and identity are clear from context, we shall denote a monoid simply by its set $M$. Given, two monoids $M$ and $N$, a morphism is a function $f\colon M\rightarrow N$ such that $f(x\cdot y)=f(x)\cdot f(y)$ for all $x,y\in M$. Note that $\Sigma^{*}$ with concatenation is a monoid with $\varepsilon$ being the identity.
Given $L\subseteq\Sigma^{*}$, the syntactic congruence of $L$ in $\Sigma^{*}$ is the relation $\sim_{L}$ defined as $u\sim_{L}v$ if and only if for all $x,y\in\Sigma^{*}$, we have $xuy\in L\Leftrightarrow xvy\in L$. The set of equivalence classes in $\Sigma^{*}$ with respect to $\sim_{L}$ is denoted $\Sigma^{*}/{\sim_{L}}$. The morphism $\eta_{L}\colon\Sigma^{*}\rightarrow\Sigma^{*}/{\sim_{L}}$ which maps words to their equivalence class is called the syntactic morphism of $L$.
Given a language $L\subseteq\Sigma^{*}$, the set $\Sigma^{*}/{\sim_{L}}$ with the multiplication $\eta_{L}(u)\cdot\eta_{L}(v)=\eta_{L}(uv)$ forms a monoid called the syntactic monoid of $L$, which we often denote with $M_{L}$, see Pin [15]. We say that $x\in M_{L}$ is aperiodic if there exists some $n\in\mathbb{N}_{+}$ such that $x^{n}=x^{n+1}$; otherwise, $x$ is periodic. A monoid $M$ is aperiodic if every element is aperiodic (see Schützenberger [17]). We note that for any finite monoid $M$, and any $x\in M$, there exist $i,p>0$ such that $x^{pn+i}=x^{i}$ for all $n\in\mathbb{N}$ (e.g., see Section 6, Chapter 2 of Pin [15]). We call an element $i\in M$ of a monoid idempotent if $i=i\cdot i$.
For a monoid $M$, a subset $G\subseteq M$ is a subgroup if $G$ forms a group (with the same multiplication as $M$, but with a potentially different identity element than $M$), and if $|G|=1$ then $G$ is a trivial subgroup.
The Logic FC
We assume the reader is familiar with the standard concepts of first-order logic (for example, see Libkin [20]). However, we shall look at the particular logic with which this paper is concerned in more detail. The definitions given here are based on the definitions given by Thompson and Freydenberger in [11] – who define $\mathsf{FC}$ in a slightly more technical way than its original definition by Freydenberger and Peterfreund [1].
$\mathsf{FC}$ is built on one fixed signature $\tau_{\Sigma}:=\{R_{\circ},\varepsilon,\{\mathtt{a}\}_{\mathtt{a}\in\Sigma}\}$ for every terminal alphabet $\Sigma$, where $R_{\circ}$ is a ternary relation symbol and where $\varepsilon$ and each $\mathtt{a}\in\Sigma$ is a constant symbol. Given a word $w\in\Sigma^{*}$, the $\tau_{\Sigma}$-structure that represents $w\in\Sigma^{*}$ is defined as $\mathcal{A}_{w}:=(A,R_{\circ}^{\mathcal{A}_{w}},\varepsilon^{\mathcal{A}_{w}},\{\mathtt{a}^{\mathcal{A}_{w}}\}_{\mathtt{a}\in\Sigma})$ where
  • $A:=\mathsf{facts}(w)\mathrel{\cup}\{\perp\}$ is the universe,
  • $R_{\circ}^{\mathcal{A}_{w}}:=\{(a,b,c)\in\mathsf{facts}(w)^{3}\mid a=b\cdot c\}$
  • $\mathtt{a}^{\mathcal{A}_{w}}:=\mathtt{a}$ if $|w|_{\mathtt{a}}\geqslant 1$, and $\mathtt{a}^{\mathcal{A}_{w}}=\perp$ otherwise, and
  • $\varepsilon^{\mathcal{A}_{w}}:=\varepsilon$.
We call such a structure $\mathcal{A}_{w}$ an $\mathsf{FC}$-structure.
Note that if $|w|_{\mathtt{a}}=0$, then $\mathtt{a}^{\mathcal{A}_{w}}=\perp$. However, we usually deal with those words where $|w|_{\mathtt{a}}\geqslant 1$ for all $\mathtt{a}\in\Sigma$. Therefore, we tend to write $\mathtt{a}\in\Sigma$ rather than $\mathtt{a}^{\mathcal{A}_{w}}\in A$.
Let $X$ be a fixed, countably infinite set of variables (where $X$ is disjoint from $\Sigma$ and $\tau_{\Sigma}$). An $\mathsf{FC}$-formula is a first-order formula where the atomic formulas are of the form $R_{\circ}(x,y,z)$, where $x$, $y$, and $z$ are variables or constants. As syntactic sugar, we write $(x\mathbin{\dot{=}}y\cdot z)$ for atomic $\mathsf{FC}$ formulas, as we always interpret $R_{\circ}$ as concatenation. More formally:
Definition 1.
Let $\mathsf{FC}$ be the set of all $\mathsf{FC}$-formulas defined recursively as:
  • If $x,y,z\in X\mathrel{\cup}\Sigma\mathrel{\cup}\{\varepsilon\}$, then $(x\mathbin{\dot{=}}y\cdot z)\in\mathsf{FC}$,
  • if $\varphi,\psi\in\mathsf{FC}$, then $(\varphi\land\psi),(\varphi\lor\psi),\neg\varphi\in\mathsf{FC}$, and
  • if $\varphi\in\mathsf{FC}$ and $x\in X$, then $\forall x\colon\varphi\in\mathsf{FC}$ and $\exists x\colon\varphi\in\mathsf{FC}$.
We allow atomic formulas of the form $x\mathbin{\dot{=}}y$, as this can be expressed by $x\mathbin{\dot{=}}y\cdot\varepsilon$. We freely omit parentheses when the meaning is clear. We also use $Qx_{1},x_{2},\dots,x_{n}\colon\varphi$ as shorthand for $Qx_{1}\colon Qx_{2}\colon\dots Qx_{n}\colon\varphi$ where $Q\in\{\exists,\forall\}$.
In $\mathsf{FC}$, an interpretation $\mathcal{I}:=(\mathcal{A}_{w},\sigma)$ consists of a $\tau_{\Sigma}$-structure $\mathcal{A}_{w}$ that represents some $w\in\Sigma^{*}$, and a mapping $\sigma\colon X\rightarrow\mathsf{facts}(w)$. Notice that $\sigma(x)\neq\perp$ is assumed for all $x\in X$. We write $\mathcal{I}\models\varphi$ to denote that $\varphi$ is true in $\mathcal{I}$, defined in the usual way (see Chapter 2 of [20] for example).
If $\varphi$ is a sentence (that is, $\varphi$ has no free variables), then we simply write $\mathcal{A}_{w}\models\varphi$. Furthermore, as $\mathcal{A}_{w}$ and $\mathcal{A}_{v}$ are isomorphic if and only if $w=v$, we can use $w$ as a shorthand for $\mathcal{A}_{w}$ when appropriate.
Definition 2.
The language defined by a sentence $\varphi\in\mathsf{FC}$ is $\mathcal{L}(\varphi):=\{w\in\Sigma^{*}\mid w\models\varphi\}$. Let $\mathcal{L}(\mathsf{FC})$ be the class of languages definable by an $\mathsf{FC}$ sentence.
In contrast to $\mathsf{FO}[<]$, $\mathsf{FC}$ can define non-regular languages. Possibly the most straightforward example is $\exists\,\mathsf{w},x\colon(\mathsf{w}\mathbin{\dot{=}}x\cdot x)$ where, as shorthand, $\mathsf{w}$ represents the whole input word (this can be easily expressed in $\mathsf{FC}$, see Example 2.4 of [11]). Now consider the formula $\varphi:=\exists\,\mathsf{w},x\colon(\mathsf{w}\mathbin{\dot{=}}x\cdot x)\land\neg\exists y\colon(y\mathbin{\dot{=}}\mathtt{b})$. Assuming $\Sigma=\{\mathtt{a},\mathtt{b}\}$, we have that $\mathcal{L}(\varphi)=(\mathtt{aa})^{*}$ which is a regular language. However, it cannot be expressed in $\mathsf{FO}[<]$ as it is not star-free; e.g., see [21].
III Main Results
In this section, we provide precise statements of this paper’s main results: A decidable characterization of the $\mathsf{FC}$-definable regular languages, including a regular expression-based characterization, an algebraic characterization, and an automata characterization. Before formulating these main results, we first need to provide the necessary definitions.
III-A Star-Free Closure
Familiarity of the basics regarding regular languages is assumed (for example, see Pin [22]). The star-free closure is an operator on classes of languages that is of particular interest to this article. Place and Zeitoun [14] conducted a systematic study of this operator.
Definition 3 (Star-free closure [14]).
Let $\mathcal{C}$ be an arbitrary class of regular languages. The class $\mathsf{SF}(\mathcal{C})$ is defined as the smallest class of languages containing $\mathcal{C}$, singletons $\{\mathtt{a}\}$ where $\mathtt{a}\in\Sigma$, and which is closed under union, complement, and concatenation.
An observation made in [14] is that the star-free closure of any class containing the empty set and $\Sigma^{*}$ contains the star-free languages. Thus, the star-free closure naturally extends the star-free languages (the regular languages definable in $\mathsf{FO}[<]$); see [14, 17, 2] for more details.
This paper is particularly interested in the star-free closure of a particular class: Let $\mathcal{R}$ denote the class of languages of the form $\{w^{*}\mid w\in\Sigma^{*}\}$. Recalling Definition 3, the class $\mathsf{SF}(\mathcal{R})$ is recursively defined as follows:
  • $\{\mathtt{a}\}\in\mathsf{SF}(\mathcal{R})$ for all $\mathtt{a}\in\Sigma$ (we often write $\mathtt{a}$ for $\{\mathtt{a}\}$), and $w^{*}\in\mathsf{SF}(\mathcal{R})$ for any $w\in\Sigma^{*}$.
  • If $L,L^{\prime}\in\mathsf{SF}(\mathcal{R})$, then each of the languages  $L\mathrel{\cup}L^{\prime}$,  $\Sigma^{*}\setminus L$,  $L\cdot L^{\prime}$  belong to $\mathsf{SF}(\mathcal{R})$.
Note that $\mathsf{SF}(\mathcal{R})$ is closed under Boolean operations (union, intersection, complement/set difference). Furthermore, notice that $\emptyset,\Sigma^{*},\{\varepsilon\}\in\mathsf{SF}(\mathcal{R})$ since $\emptyset=(\mathtt{a}\mathrel{\cap}\mathtt{b})$, $\Sigma^{*}=\Sigma^{*}\setminus\emptyset$, and $\{\varepsilon\}=\mathtt{a}^{*}\setminus\mathtt{a}\cdot\mathtt{a}^{*}$ for $\mathtt{a},\mathtt{b}\in\Sigma$ with $\mathtt{a}\neq\mathtt{b}$.
Example 4.
Let $\Sigma=\{\mathtt{a},\mathtt{b}\}$ and consider the regular language $L:=(\mathtt{aa}\mathrel{\cup}\mathtt{bb})^{*}$. While it may not be immediately clear that $L$ belongs to $\mathsf{SF}(\mathcal{R})$, consider the following $\mathsf{SF}(\mathcal{R})$-language:
$L_{1}:=\bigl{(}(\{\varepsilon\}\mathrel{\cup}\Sigma^{*}\mathtt{b})\cdot(\mathtt{aa})^{*}\mathtt{a}\cdot(\mathtt{b}\Sigma^{*}\mathrel{\cup}\{\varepsilon\})\bigr{)}\mathrel{\cup}\\ \bigl{(}(\{\varepsilon\}\mathrel{\cup}\Sigma^{*}\mathtt{a})\cdot(\mathtt{bb})^{*}\mathtt{b}\cdot(\mathtt{a}\Sigma^{*}\mathrel{\cup}\{\varepsilon\})\bigr{)}.$
For any $w\in L$ and any factor $\mathtt{a}^{n}\sqsubseteq w$ which is a prefix or preceded by $\mathtt{b}$, and is a suffix or succeeded by $\mathtt{b}$, it necessarily holds that $n$ is even. The symmetric holds for the analogous factors $\mathtt{b}^{n}\sqsubseteq w$. One can therefore verify that $\Sigma^{*}\setminus L_{1}=L$, and thus $L$ does indeed belong to $\mathsf{SF}(\mathcal{R})$.
III-B Group Primitive Languages
We now introduce a new class of regular languages defined by a restriction on the syntactic morphism:
Definition 5 (Group Primitive Language).
Let $M$ be a finite monoid. A morphism $\mu\colon\Sigma^{*}\rightarrow M$ is group primitive if $|\varrho(\mu^{-1}(x))|=1$ for all periodic elements $x\in M$. A language $L\subseteq\Sigma^{*}$ is group primitive if it is regular and its syntactic morphism is group primitive.
Let us note the connection to languages recognized by aperiodic monoids. Recall that an aperiodic monoid is a monoid where all elements $x$ satisfy $x^{n}=x^{n+1}$ for some $n\in\mathbb{N}_{+}$. This can alternatively be phrased as follows: An aperiodic monoid is a monoid where all subgroups are trivial (see [17], for example). In comparison, while the syntactic monoid of a group primitive language may contain non-trivial subgroups, the preimage of any element of a non-trivial subgroup (under the syntactic morphism) is a language of the form $\{w^{i}\mid i\in I\}$ for some $w\in\Sigma^{*}$ and $I\subseteq\mathbb{N}$. This is because if $G\subseteq M_{L}$ is a subgroup, then the identity of $G$ is the only aperiodic element (if we take $x^{n}=x^{n+1}$ with $x\in G$, then a simple cancellation argument implies $x$ is the identity of $G$).
Example 6.
Let $\Sigma=\{\mathtt{a},\mathtt{b}\}$. The language $L_{1}:=(\mathtt{aa})^{*}$ has the syntactic monoid $\{e,x,y\}$ with $e$ being the identity, $xx=e$, and $yx=xy=yy=y$. The syntactic morphism $\eta_{L_{1}}\colon\Sigma^{*}\rightarrow\{e,x,y\}$ is: $\eta_{L_{1}}(w)=e$ if $w\in(\mathtt{aa})^{*}$, $\eta_{L_{1}}(w)=x$ if $w\in\mathtt{a}(\mathtt{aa})^{*}$, and $\eta_{L_{1}}(w)=y$ if $|w|_{\mathtt{b}}\geqslant 1$. Clearly, $x$ is the only periodic element, however as $\eta_{L_{1}}^{-1}(x)=\mathtt{a}(\mathtt{aa})^{*}$, we have that $L_{1}$ is a group primitive language since $\varrho(\mathtt{a}(\mathtt{aa})^{*})=\{\mathtt{a}\}$.
Now consider $L_{2}:=\{w\in\Sigma^{*}\mid|w|_{\mathtt{a}}\text{ is even}\}$. The syntactic monoid for $L_{2}$ is $\{e,x\}$ with $e$ being the identity and $xx=e$. The syntactic morphism $\eta_{L_{2}}\colon\Sigma^{*}\rightarrow\{e,x\}$ is $\eta_{L_{2}}(w)=e$ if $|w|_{\mathtt{a}}$ is even, and $\eta_{L_{2}}(w)=x$ otherwise. Again, it is clear that $x$ is periodic, however, $\mathtt{ba}$ and $\mathtt{baaa}$ are in $\eta_{L_{2}}^{-1}(x)$ which means that $L_{2}$ is not group primitive as $\varrho(\mathtt{ba})\neq\varrho(\mathtt{baaa})$.
III-C Automata with Loop-Step Cycles
We denote a deterministic finite automaton (DFA) $\mathcal{M}$ by $(Q,\Sigma,\delta,q_{0},F)$ with $Q$ being the set of states, $\Sigma$ being the alphabet, $\delta\colon Q\times\Sigma\rightarrow Q$ being the transition function, $q_{0}\in Q$ being the start state, and $F\subseteq Q$ being the set of accepting states. We write $\delta^{*}\colon Q\times\Sigma^{*}\rightarrow Q$ as the reflexive and transitive closure of $\delta$. Then, $\mathcal{M}$ defines the language $\mathcal{L}(\mathcal{M})$ of all words $w\in\Sigma^{*}$ such that $\delta^{*}(q_{0},w)\in F$. We call a DFA minimal if there does not exist an equivalent DFA with fewer states.
We now define a condition for minimal DFAs:
Definition 7 (Loop-Step Cycle).
Let $\mathcal{M}:=(Q,\Sigma,\delta,q_{0},F)$ be a minimal DFA. We say that $\mathcal{M}$ has a loop-step cycle if there exist $n\geqslant 2$ pairwise distinct states $p_{0},p_{1},\dots,p_{n-1}$ and words $w,v\in\Sigma^{+}$, where $\varrho(w)\neq\varrho(v)$, such that:
  • $\delta^{*}(p_{i},w)=p_{i}$ for all $i\in\{0,\dots,n-1\}$, and
  • $\delta^{*}(p_{i},v)=p_{i+1}$ for all $i\in\{0,\dots,n-2\}$ and $\delta^{*}(p_{n-1},v)=p_{0}$.
We shall often write $\delta^{*}(p_{i},v)=p_{i+1\pmod{n}}$ to denote that $\delta^{*}(p_{i},v)=p_{i+1}$ for all $i\in\{0,\dots,n-2\}$ and that $\delta^{*}(p_{n-1},v)=p_{0}$.
Example 8.
Consider the languages $L_{1}:=(\mathtt{aa}\mathrel{\cup}\mathtt{ab}\mathrel{\cup}\mathtt{ba})^{*}$ and $L_{2}:=(\mathtt{aa}\mathrel{\cup}\mathtt{ab}\mathrel{\cup}\mathtt{bb})^{*}$. The minimal DFA for $L_{1}$ is given in Fig. 2 (on the left-hand side) and the minimal DFA for $L_{2}$ is given in Fig. 2 (on the right-hand side). Note that the automaton for $L_{1}$ has a loop-step cycle since $\delta(p_{0},\mathtt{a})=p_{1}$ and $\delta(p_{1},\mathtt{a})=p_{0}$, and $\delta^{*}(\mathtt{ba},p_{0})=p_{0}$ and $\delta^{*}(\mathtt{ba},p_{1})=p_{1}$. Although it is somewhat tedious to verify, the automaton for $L_{2}$ does not have a loop-step cycle: To see why this holds, assume $\mathcal{M}_{2}$ is the automaton given in Fig. 2 for $L_{2}$, and $\mathcal{M}_{2}$ has a loop-step cycle. Hence, there are $n\geqslant 2$ pairwise distinct states $p_{0},p_{1},\dots,p_{n-1}$ and $w,v\in\Sigma^{*}$ where $\varrho(w)\neq\varrho(v)$ such that:
  • $\delta^{*}(p_{i},w)=p_{i}$ for all $i\in\{0,\dots,n-1\}$, and
  • $\delta^{*}(p_{i},v)=p_{i+1\pmod{n}}$.
Notice that $q_{2}$ cannot reach any other state, and therefore $q_{2}\notin\{p_{0},\dots,p_{n-1}\}$. Furthermore, the only incoming transition for $q_{3}$ is labeled $\mathtt{b}$ and the only incoming transition for $q_{1}$ is labeled $\mathtt{a}$ and thus both $q_{3}$ and $q_{1}$ cannot be in $\{p_{0},\dots,p_{n-1}\}$. This leaves us with $\{p_{0},\dots,p_{n-1}\}$ being either $\{q_{0},q_{1}\}$ or $\{q_{0},q_{3}\}$. For state $q_{i}$ and $q_{j}$, let $L_{q_{i},q_{j}}:=\{u\mid\delta^{*}(q_{i},u)=q_{j}\}$. Then, one can verify that
  • $L_{q_{0},q_{0}}\mathrel{\cap}L_{q_{1},q_{1}}=(\mathtt{aa})^{*}$ and $L_{q_{0},q_{1}}\mathrel{\cap}L_{q_{1},q_{0}}=\mathtt{a}(\mathtt{aa})^{*}$ and thus $\varrho(w)=\varrho(v)$.
  • $L_{q_{0},q_{0}}\mathrel{\cap}L_{q_{3},q_{3}}=(\mathtt{bb})^{*}$ and $L_{q_{0},q_{3}}\mathrel{\cap}L_{q_{3},q_{0}}=\mathtt{b}(\mathtt{bb})^{*}$ and thus $\varrho(w)=\varrho(v)$.
Consequently, $\mathcal{M}_{2}$ cannot have a loop-step cycle.
$p_{0}$$p_{1}$$p_{2}$$p_{3}$$\mathtt{a}$$\mathtt{a},\mathtt{b}$$\mathtt{b}$$\mathtt{a}$$\mathtt{a},\mathtt{b}$$\mathtt{b}$$q_{0}$$q_{1}$$q_{2}$$q_{3}$$\mathtt{a}$$\mathtt{a},\mathtt{b}$$\mathtt{b}$$\mathtt{b}$$\mathtt{a},\mathtt{b}$$\mathtt{a}$
Figure 2: Minimal DFA for $(\mathtt{aa}\mathrel{\cup}\mathtt{ab}\mathrel{\cup}\mathtt{ba})^{*}$ on the left-hand side, and minimal DFA for $(\mathtt{aa}\mathrel{\cup}\mathtt{ab}\mathrel{\cup}\mathtt{bb})^{*}$ on the right-hand side. See Example 8.
We now show that it is decidable to determine whether an automaton has a loop-step cycle:
Theorem 9.
Deciding whether a minimal DFA has a loop-step cycle is $\mathsf{PSPACE}$-complete.
Proof:
Let $\mathcal{M}:=(Q,\Sigma,\delta,q_{0},F)$ by a minimal DFA. We shall prove that deciding whether $\mathcal{M}$ has a loop-step cycle is $\mathsf{PSPACE}$-complete.
An array $B$ of size $n$ is a cyclic shift of an array $A$ of size $n$ if $B[i]=A[i+1]$ for all $10\leqslant i<n-$ 1, and $B[n-1]=A[0]$.
Upper Bound
Upon input of a minimal DFA $\mathcal{M}$, we want to decide in $\mathsf{PSPACE}$ whether $\mathcal{M}$ has a loop-step cycle. By Savitch’s Theorem (see Section 4.3 of [23]), we have $\mathsf{PSPACE}=\mathsf{NPSPACE}$. This allows us to make non-deterministic guesses. Clearly, Algorithm 1 runs non-deterministically with polynomial space; hence it decides a problem that belongs to $\mathsf{PSPACE}$. It remains for us to show the correctness of Algorithm 1.
Algorithm 1 $\mathsf{NPSPACE}$ algorithm for the following:
Input: Minimal DFA $\mathcal{M}=(Q,\Sigma,\delta,q_{0},F)$.
Output: $\mathsf{True}$ iff $\mathcal{M}$ has a loop-step cycle.
1:Guess an integer $n$ with $2\leqslant n\leqslant|Q|$
2:Guess $n$ distinct states $p_{0},\dots,p_{n-1}\in Q$
3:Let $A$ be an array such that $A[i]\leftarrow p_{i}$ for $0\leqslant i<n$
4:Let $B$ and $C$ be arrays such that $A=B=C$
5:Let $\mathsf{diff}\leftarrow\mathsf{False}$
6:while ($B\neq C$) or ($B$ is not a cyclic shift of $A$) or ($\mathsf{diff}=\mathsf{False}$do
7:     Guess $\mathtt{a},\mathtt{b}\in\Sigma$
8:     if $\mathtt{a}\neq\mathtt{b}$ then
9:         $\mathsf{diff}\leftarrow\mathsf{True}$
10:     end if
11:     $B[i]\leftarrow\delta(B[i],\mathtt{a})$ for each $0\leqslant i<n$
12:     $C[i]\leftarrow\delta(C[i],\mathtt{b})$ for each $0\leqslant i<n$
13:end while
14:Return $\mathsf{True}$
Assume Algorithm 1 returns true. Let $u=\mathtt{a}_{1}\cdots\mathtt{a}_{m}$ be the nondeterministic guesses for $B$, and let $v=\mathtt{b}_{1}\cdots\mathtt{b}_{m}$ be the guesses for $C$. Since $\mathsf{diff}$ is true, we know that $u\neq v$. Furthermore, as $|u|=|v|$, we also know that $\varrho(u)\neq\varrho(v)$. Since $B$ and $C$ are cyclic shifts of $A$, it follows that $\delta^{*}(p_{i},u)=\delta^{*}(p_{i},v)=p_{i+1\pmod{n}}$. Hence, we have that $\delta^{*}(p_{i},u)=p_{i+1\pmod{n}}$ and for all $0\leqslant i<n$ we have $\delta^{*}(p_{i},v^{n})=p_{i}$ where $\varrho(u)\neq\varrho(v^{n})$ as $\varrho(u)\neq\varrho(v)$. This concludes one direction.
For the other direction, let $\mathcal{M}=(Q,\Sigma,\delta,q_{0},F)$ be a minimal DFA that has a loop-step cycle. Let $w,v\in\Sigma^{+}$, where $\varrho(w)\neq\varrho(v)$, and let $p_{0},\ldots,p_{n-1}$ be $n$ distinct states with:
  • $\delta^{*}(p_{i},w)=p_{i}$ for all $i\in\{0,\dots,n-1\}$, and
  • $\delta^{*}(p_{i},v)=p_{i+1\pmod{n}}$.
Let $u_{1}:=w^{n|v|}\cdot v^{n|w|+n+1}$ and $u_{2}:=w^{2n|v|}\cdot v^{n+1}$. First, we show that $u_{1}\neq u_{2}$: If $w^{n|v|}\cdot v^{n|w|+n+1}=w^{2n|v|}\cdot v^{n+1}$, then $v^{n|w|}=w^{n|v|}$ which contradicts $\varrho(w)\neq\varrho(v)$ as $n|w|\neq 0\neq n|v|$. Next, notice that $|u_{1}|=|u_{2}|$:
  • $|u_{1}|=n\cdot|w||v|+|v|(n|w|+n+1)=2n|v||w|+n|v|+|v|$,
  • $|u_{2}|=2n\cdot|w||v|+|v|(n+1)=2n|v||w|+n|v|+|v|$.
If $\mathcal{M}$ is in state $p_{i}$ for some $0\leqslant i<n$, then reading $w^{k}$ for any $k\in\mathbb{N}$ does not change the state of $\mathcal{M}$. Likewise, if $\mathcal{M}$ is in state $p_{i}$, then reading $v^{kn+1}$ for any $k\in\mathbb{N}_{+}$ puts $\mathcal{M}$ into state $p_{i+1\pmod{n}}$. Hence, for any $i\in[n]$ we have that $\delta^{*}(p_{i},u_{1})=\delta^{*}(p_{i},u_{2})=p_{i+1\pmod{n}}$.
Assume that for $j$ with $1\leqslant j\leqslant|u_{1}|$, the $j$-th time line 11 of Algorithm 1 is executed, $\mathtt{a}$ is the $j$-th letter of $u_{1}$. Likewise, assume that the $j$-th time line 12 of Algorithm 1 is executed, $\mathtt{b}$ is the $j$-th letter of $u_{2}$. Firstly, $\mathsf{diff}=\mathsf{True}$ since $u_{1}\neq u_{2}$. Moreover, since $\delta^{*}(p_{i},u_{1})=\delta^{*}(p_{i},u_{2})=p_{i+1\pmod{n}}$ we have that $B=C$ and $B$ is a cyclic shift of $A$.
Thus, Algorithm 1 returns $\mathsf{True}$.
Lower Bound
To show $\mathsf{PSPACE}$-hardness, we reduce from the $\mathsf{PSPACE}$-complete problem finite-automaton cycle existence [18] which is defined as follows: Given a minimal DFA $\mathcal{M}$, decide whether $\mathcal{M}$ has a cycle, i.e., whether there exists a word $u\in\Sigma^{*}$ and a state $p$ of $\mathcal{M}$ such that $\delta^{*}(p,u)\neq p$ and $\delta^{*}(p,u^{r})=p$ for some $r\in\mathbb{N}_{+}$.
For an instance $\mathcal{M}=(Q,\Sigma,\delta,q_{0},F)$ of the finite-automaton cycle existence problem, we construct the automaton $\mathcal{M}_{2}=(Q_{2},\Sigma_{2},\delta_{2},q_{0,2},F_{2})$ where
  • $Q_{2}:=Q$,
  • $\Sigma_{2}:=\Sigma\mathrel{\cup}\{\bar{\mathtt{a}}\}$, where $\bar{\mathtt{a}}$ is a new letter with $\bar{\mathtt{a}}\not\in\Sigma$,
  • for all $q\in Q$, we have $\delta_{2}(q,\bar{\mathtt{a}})=q$, and for any $q\in Q$ and $\mathtt{a}\in\Sigma$, we have $\delta_{2}(q,\mathtt{a}):=\delta(q,\mathtt{a})$,
  • $q_{0,2}:=q_{0}$, and
  • $F_{2}:=F$.
Since $\mathcal{M}$ is a minimal DFA, $\mathcal{M}_{2}$ is also a minimal DFA. Clearly, constructing $\mathcal{M}_{2}$ from $\mathcal{M}$ can be done in polynomial time with respect to the size of $\mathcal{M}$. We now show that $\mathcal{M}_{2}$ has a loop-step cycle $\iff \mathcal{M}$ has a cycle.
[…]
For direction $\Longrightarrow$” assume that $\mathcal{M}$ does not have a cycle, i.e., there does not exist $u\in\Sigma^{*}$ and a state $q\in Q$ such that $\delta^{*}(q,u)\neq q$ and $\delta^{*}(q,u^{r})=q$ for some $r\in\mathbb{N}_{+}$. Since $\mathcal{M}_{2}$ is obtained from $\mathcal{M}$ by only adding self-loops $\delta_{2}(p,\bar{\mathtt{a}})=p$ for each state $p$, it necessarily holds that there does not exist $u^{\prime}\in\Sigma_{2}^{*}$ and a state $q^{\prime}\in Q_{2}$ such that $\delta_{2}^{*}(q^{\prime},u^{\prime})\neq q^{\prime}$ and $\delta_{2}^{*}(q^{\prime},{u^{\prime}}^{r})=q^{\prime}$ for some $r\in\mathbb{N}_{+}$. Thus, $\mathcal{M}_{2}$ cannot have a loop-step cycle.
In summary, we have provided a polynomial-time reduction from the $\mathsf{PSPACE}$-complete problem finite-automaton cycle existence to the problem of deciding whether a given minimal DFA has a loop-step cycle. This shows that the latter problem is $\mathsf{PSPACE}$-hard.
III-D Main Theorems
With the prerequisites out of the way, we are now ready to state our main result:
Theorem 10.
Let $L\subseteq\Sigma^{*}$ be a regular language, and let $\mathcal{M}:=(Q,\Sigma,\delta,q_{0},F)$ be a minimal DFA with $L=\mathcal{L}(\mathcal{M})$. Then, the following are equivalent:
  1. 1.
    $L\in\mathcal{L}(\mathsf{FC})$,
  2. 2.
    $L\in\mathsf{SF}(\mathcal{R})$,
  3. 3.
    $L$ is group primitive,
  4. 4.
    $\mathcal{M}$ does not have a loop-step cycle.
This provides a characterization of the regular languages that can be expressed in $\mathsf{FC}$. Our second main result is that this characterization is decidable, in fact (for minimal automata) it is $\mathsf{PSPACE}$-complete.
Theorem 11.
Given a minimal DFA $\mathcal{M}$, deciding whether $\mathcal{L}(\mathcal{M})\in\mathcal{L}(\mathsf{FC})$ is $\mathsf{PSPACE}$-complete.
Theorem 11 is immediately obtained by combining Theorem 10 with Theorem 9.
With the main characterization given (Theorem 10) and the result that this characterization is decidable (Theorem 11), the rest of this article is devoted to proving Theorem 10:
[…]
IV On Loop-Step Cycles and Group Primitive Regular Languages
In this section we prove the following theorem.
Theorem 12.
Let $L$ be a regular language and let $\mathcal{M}$ be a minimal DFA with $L=\mathcal{L}(\mathcal{M})$. If $L$ is not group primitive, then $\mathcal{M}$ has a loop-step cycle.
Proof:
Let $L$ be a regular language that is not group primitive. Let $\mathcal{M}:=(Q,\Sigma,\delta,q_{0},F)$ be a minimal DFA with $L=\mathcal{L}(\mathcal{M})$. Let $\eta_{L}\colon\Sigma^{*}\rightarrow M_{L}$ be the syntactic morphism of $L$, where $M_{L}$ is the syntactic monoid for $L$. Since $L$ is not group primitive, $\eta_{L}$ is not group primitive. I.e., there exists a periodic element $g\in M_{L}$ where $|\varrho(\eta_{L}^{-1}(g))|\geqslant 2$. Hence, there exist words $w,v\in\eta_{L}^{-1}(g)$ such that $\varrho(w)\neq\varrho(v)$.
Since $g$ is periodic, we have that $g^{n}\neq g^{n+1}$ for any $n\in\mathbb{N}$. However, there exists $y\geqslant 0$ and $m>0$ such that for all $x\in\mathbb{N}_{+}$, we get $g^{y+mx}=g^{y+m}$; for example, see Section 6, Chapter 2 of [15]. We choose $m$ as small as possible. Note that $m\geqslant 2$ because $g$ is periodic. We let $G:=\{g^{y+m},g^{y+m+1},\dots,g^{y+2m-1}\}$. Observe that $G$ has $m$ elements, and $G$ is closed under multiplication with $g$, i.e., $g\cdot h=h\cdot g\in G$ for every $h\in G$.
For any words $u,u^{\prime}\in\Sigma^{*}$ we have: $\eta_{L}(u)=\eta_{L}(u^{\prime})$ if and only if $\delta^{*}(q,u)=\delta^{*}(q,u^{\prime})$ for all $q\in Q$; see Proposition 4.28 of [15]. By our choice of $w$ and $v$ we have: $w,v\in\Sigma^{+}$, $\varrho(w)\neq\varrho(v)$, and $\eta_{L}(w)=\eta_{L}(v)=g$.
In particular, for any $z,z^{\prime}\in\mathbb{N}$ with $z,z^{\prime}\geqslant y+m$ we have: $v^{z}\sim_{L}v^{z^{\prime}}$ $\iff$ $\eta_{L}(v^{z})=\eta_{L}(v^{z^{\prime}})$ $\iff$ $g^{z}=g^{z^{\prime}}$ $\iff$ $z-y\equiv z^{\prime}-y\pmod{m}$. Thus, since $m\not\equiv m{+}1\pmod{m}$, we have: $v^{m+y}\not\sim_{L}v^{m+y+1}$. Hence, there exists some state $p\in Q$ such that $\delta^{*}(p,v^{m+y})\neq\delta^{*}(p,v^{m+y+1})$. For every $i\in\{0,\ldots,m{-}1\}$ let $p_{i}:=\delta^{*}(p,v^{m+y+i})$. Note that $p_{0}\neq p_{1}$ (because $p_{0}=\delta^{*}(p,v^{m+y})\neq\delta^{*}(p,v^{m+y+1})=p_{1}$), and that $\delta^{*}(p_{i},v)=p_{i+1}$ holds for every $i\in\{0,\ldots,m-2\}$.
Furthermore, $\delta^{*}(p_{m-1},v)=p_{0}$ due to the following reasoning: Obviously, $\delta^{*}(p_{m-1},v)=\delta^{*}(p,v^{y+2m})$. Furthermore, $v^{y+2m}\sim_{L}v^{y+m}$ (since $\eta_{L}(v^{y+2m})=g^{y+2m}=g^{y+m}=\eta_{L}(v^{y+m})$). Thus, for every state $q\in Q$ we have $\delta^{*}(q,v^{y+2m})=\delta^{*}(q,v^{y+m})$. In particular for $q=p$ this yields: $\delta^{*}(p,v^{y+2m})=\delta^{*}(p,v^{y+m})=p_{0}$.
In summary, we now know that $\delta^{*}(p_{i},v)=p_{i+1\pmod{m}}$. But note that the states $p_{0},p_{1},\dots,p_{m-1}$ are not necessarily pairwise distinct. However, we know that, for all $x\in\mathbb{N}_{+}$, we have that $v^{m+y}\sim_{L}v^{mx+y}$. Therefore, we get a set of states $\{q_{0},q_{1},\dots,q_{m^{\prime}-1}\}$ of size $m^{\prime}$, where $2\leqslant m^{\prime}\leqslant m$, such that $q_{0}=p_{0}$, $q_{1}=p_{1}$ and $\delta^{*}(q_{i},v)=q_{i+1\pmod{m^{\prime}}}$. This is because we know that we eventually need to “loop back” to state $p_{0}$, which prohibits $\delta^{*}(p_{i},v)=p_{j}$ where $p_{j}\neq p_{0}\neq p_{i}$, and thus we obtain our set $\{q_{0},q_{1},\dots,q_{m^{\prime}-1}\}$.
Since $v^{m+y}\sim_{L}v^{mx+y}$ for all $x\in\mathbb{N}_{+}$, it follows that $\delta^{*}(q,v^{m+y})=\delta^{*}(q,v^{mx+y})$ for all $x\in\mathbb{N}_{+}$ and all $q\in Q$. Furthermore, we have that $w,v\in\eta_{L}^{-1}(g)$, thus $w^{z}\sim_{L}v^{z}$ for all $z\in\mathbb{N}$. In particular, this implies for all $i\in\{0,\ldots,m^{\prime}{-}1\}$ and the state $q_{i}$ that $\delta^{*}(q_{i},w^{m+y})=q_{i}$ for all $0\leqslant i<m^{\prime}$. We let $w^{\prime}:=w^{m+y}$ and obtain that $\delta^{*}(q_{i},w^{\prime})=q_{i}$ for all $i\in\{0,\ldots,m^{\prime}{-}1\}$. Furthermore, $\varrho(w^{\prime})=\varrho(w)\neq\varrho(v)$.
In summary, we now know that $q_{0},\ldots,q_{m^{\prime}-1}$ are $m^{\prime}\geqslant 2$ pairwise distinct states, $w^{\prime}$ and $v$ are words with $\varrho(w^{\prime})\neq\varrho(v)$, and $\delta^{*}(q_{i},w^{\prime})=q_{i}$ and $\delta^{*}(q_{i},v)=q_{i+1\pmod{m}}$ for all $i$ where $10\leqslant i<m^{\prime}{-}$ 1. This proves that $\mathcal{M}$ has a loop-step cycle and therefore completes the proof of Theorem 12.
[…]
V Every $\mathsf{SF}(\mathcal{R})$-Language is Expressible in $\mathsf{FC}$
In this section, we show that the star-free closure of the class $\{w^{*}\mid w\in\Sigma^{*}\}$ can be expressed in $\mathsf{FC}$. We know from Example 3.7 of [1] that $\mathsf{FC}$ can express all the star-free languages, and moreover, Lemma 5.5 of [1] shows that $w^{*}$ for any $w\in\Sigma^{*}$ can be expressed in $\mathsf{FC}$. It is therefore unsurprising that we can combine these results to show $\mathsf{SF}(\mathcal{R})\subseteq\mathcal{L}(\mathsf{FC})$.
A language $L\subseteq\Sigma^{*}$ is bounded if $L\subseteq w_{1}^{*}\cdots w_{n}^{*}$ for some $w_{1},\dots,w_{n}\in\Sigma^{+}$ and $n\geqslant 1$. A language is a bounded regular language if it is both bounded and regular (Ginsburg and Spanier [25]). Lemma 5.3 of [11] states that the Boolean combination of bounded regular languages can be expressed in $\mathsf{FC}$. Going only slightly beyond that result gives us:
Theorem 13.
If $L\in\mathsf{SF}(\mathcal{R})$, then $L\in\mathcal{L}(\mathsf{FC})$.
Proof:
This proof proceeds with a straightforward structural induction. Perhaps the only non-obvious step is showing $w^{*}\in\mathcal{L}(\mathsf{FC})$ for any $w\in\Sigma^{*}$. For this, we use a result on commuting words: If $uv=vu$ with $u,v\in\Sigma^{*}$, then there exists $n,m\in\mathbb{N}$ and $w\in\Sigma^{*}$ with $u=w^{n}$ and $v=w^{m}$ (See Proposition 1.3.2 of Lothaire [26]). The use of this result to show $w^{*}\in\mathcal{L}(\mathsf{FC})$ is not new to this paper: For example, see the proof of Lemma 5.5 of [1] given in [27].
[…]
VI Every Group Primitive Language is in $\mathsf{SF}(\mathcal{R})$
In this section, we show that any group primitive language (Definition 5) belongs to $\mathsf{SF}(\mathcal{R})$ (recall Section III-A). To show this, we consider many concepts and results from Place and Zeitoun [14]; which algebraically characterize the star-free closure of so-called prevarieties. We shall consider a prevariety $\mathcal{B}q$ (studied in Daviaud and Paperman [28]) such that $\mathsf{SF}(\mathcal{B}q)=\mathsf{SF}(\mathcal{R})$, and then show that all group primitive languages satisfy the algebraic characterization of $\mathsf{SF}(\mathcal{B}q)$ which follows from [14]. Consequently, all group primitive languages belong to $\mathsf{SF}(\mathcal{R})$.
For $L\subseteq\Sigma^{*}$ and $u\in\Sigma^{*}$, let $u^{-1}L:=\{w\in\Sigma^{*}\mid uw\in L\}$ and let $Lu^{-1}:=\{w\in\Sigma^{*}\mid wu\in L\}$. A class $\mathcal{C}$ of languages is a prevariety if $\mathcal{C}$ is a subset of the regular languages with $\emptyset,\Sigma^{*}\in\mathcal{C}$, for any $L\in\mathcal{C}$ and $u\in\Sigma^{*}$ we have $Lu^{-1},u^{-1}L\in\mathcal{C}$, and $\mathcal{C}$ is closed under union and complement.
Example 14.
Consider the star-free languages, which we denote with $\mathsf{SF}$. This class is the smallest class that contains the finite languages, and is closed under union, concatenation and complement. Clearly, $\mathsf{SF}$ is contained in the regular languages, and it is straightforward to show that $\emptyset$ and $\Sigma^{*}$ belong to this class. Furthermore, $\mathsf{SF}$ is closed under union and complement by definition. Thus, $\mathsf{SF}$ is a prevariety since $Lu^{-1},u^{-1}L\in\mathsf{SF}$ for all $L\in\mathsf{SF}$ and $u\in\Sigma^{*}$; for example, see Pin [21].
[…]
Definition 15 (Adapted from [14]).
Let $\mathcal{C}$ be a prevariety. If $\mu\colon\Sigma^{*}\rightarrow M$ is a morphism where $M$ is a finite monoid, then for any idempotent element $i\in M$, the $\mathcal{C}$-orbit of $i$ with respect to $\mu$ is defined as the set of all elements $isi\in M$ where $s\in M$ such that $\mu^{-1}(s)$ is not $\mathcal{C}$-separable from $\mu^{-1}(i)$.
Note that $\mu^{-1}(s)$ is not $\mathcal{C}$-separable from $\mu^{-1}(i)$ if and only if $\mu^{-1}(s)\mathrel{\cap}L\neq\emptyset$ for any $L\in\mathcal{C}$ with $\mu^{-1}(i)\subseteq L$.
For a morphism $\mu\colon\Sigma^{*}\rightarrow M$ where $M$ is finite, every idempotent element $i\in M$ defines a subset of $M$ called the $\mathcal{C}$-orbit of $i$ with respect to $\mu$. Lemma 5.5 of [14] states that for a prevariety $\mathcal{C}$, any $\mathcal{C}$-orbit is a monoid. Furthermore, every morphism $\mu\colon\Sigma^{*}\rightarrow M$ gives rise to a set of monoids called the $\mathcal{C}$-orbits of $\mu$; where every idempotent $i\in M$ is associated to some monoid in this set. Using this notion of $\mathcal{C}$-orbits, Place and Zeitoun [14] algebraically characterized the languages belonging to $\mathsf{SF}(\mathcal{C})$, whenever $\mathcal{C}$ is a prevariety:
Theorem 16 (Theorem 5.11, [14]).
Let $\mathcal{C}$ be a prevariety and let $L\subseteq\Sigma^{*}$ be regular. Then $L\in\mathsf{SF}(\mathcal{C})$ if and only if all the $\mathcal{C}$-orbits of the syntactic morphism are aperiodic monoids.
In order to use Theorem 16, we define the prevariety $\mathcal{B}q$ (introduced in [28]) as follows:
  • $\emptyset,\Sigma^{*}\in\mathcal{B}q$,
  • $L\in\mathcal{B}q$ for every regular language $L\subseteq w^{*}$ with $w\in\Sigma^{*}$,
  • $\mathcal{B}q$ is closed under union and complement, and
  • $u^{-1}L,Lu^{-1}\in\mathcal{B}q$ for any $L\in\mathcal{B}q$ and $u\in\Sigma^{*}$.
Even though $\mathcal{R}$ is not a prevariety, the following lemma establishes that $\mathsf{SF}(\mathcal{R})$ can be characterized with $\mathcal{C}$-orbits.
Lemma 17.
$\mathsf{SF}(\mathcal{R})=\mathsf{SF}(\mathcal{B}q)$.
Proof:
The $\mathsf{SF}(\mathcal{R})\subseteq\mathsf{SF}(\mathcal{B}q)$ direction is immediate since $\mathcal{R}\subseteq\mathcal{B}q$ by definition. For the other direction, it suffices to show that for any $L\in\mathcal{B}q$, we have that $L\in\mathsf{SF}(\mathcal{R})$. Thus, the proof proceeds with a straightforward structural induction along the recursive definition of $\mathcal{B}q$. Some effort is needed to deal with $u^{-1}L$; however, this is handled using a sub-induction along the length of $u$.
We note that Theorem 16 and Lemma 17 immediately give us an algebraic characterization of $\mathsf{SF}(\mathcal{R})$. However, for our more specialized purposes, we are able to give a more lightweight characterization.
Lemma 18.
If the syntactic morphism $\eta_{L}\colon\Sigma^{*}\rightarrow M_{L}$ of the regular language $L\subseteq\Sigma^{*}$ is a group primitive morphism, then all $\mathcal{B}q$-orbits of $\eta_{L}$ are aperiodic monoids.
Proof:
Let $\eta_{L}\colon\Sigma^{*}\rightarrow M_{L}$ be the syntactic morphism of the regular language $L\subseteq\Sigma^{*}$. Working towards a contradiction, assume that $\eta_{L}$ is group primitive, and for some idempotent element $i\in M_{L}$, the $\mathcal{B}q$-orbit of $i$ with respect to $\eta_{L}$ contains some periodic element $isi\in M_{L}$; that is, $(isi)^{n}\neq(isi)^{n+1}$ for all $n\in\mathbb{N}_{+}$. We fix such $i,s\in M_{L}$ for the remainder of this proof. Since $i\in M_{L}$ is idempotent:
$(isi)^{n}=\underbrace{isi\cdot isi\cdots isi}_{n\text{-times}}=\underbrace{is\cdot is\cdots is}_{n\text{-times}}\cdot i=(is)^{n}i$
for $n\in\mathbb{N}_{+}$. Furthermore, since $(isi)^{n}\neq(isi)^{n+1}$ we know that $(is)^{n}i\neq(is)^{n+1}i$ for all $n\in\mathbb{N}_{+}$. Therefore $is\in M_{L}$ is a periodic element as $(is)^{n}\neq(is)^{n+1}$ for all $n\in\mathbb{N}_{+}$.
Before moving on, we show that $\varrho(\eta_{L}^{-1}(s))\neq\emptyset$ and $\varrho(\eta_{L}^{-1}(i))\neq\emptyset$ necessarily holds. In other words, both $\eta_{L}^{-1}(s)$ and $\eta_{L}^{-1}(i)$ contain a non-empty word. First, working towards a contradiction, assume that $\varrho(\eta_{L}^{-1}(s))=\emptyset$. Then, $\eta_{L}^{-1}(s)=\{\varepsilon\}$ holds. This implies that $s$ is the identity element of $M_{L}$ and hence $(is)^{n}=i^{n}=i^{n+1}=(is)^{n+1}$ which contradicts our assumption that $is$ is periodic. Next, working towards a contradiction, assume that $\varrho(\eta_{L}^{-1}(i))=\emptyset$ and hence $\eta_{L}^{-1}(i)=\{\varepsilon\}$. Since $\{\varepsilon\}\in\mathcal{B}q$ by considering $\mathtt{a}^{*}\mathrel{\cap}\mathtt{b}^{*}$ with $\mathtt{a}\neq\mathtt{b}$, we know that $\eta_{L}^{-1}(s)\mathrel{\cap}\{\varepsilon\}\neq\emptyset$ by the definition of a $\mathcal{B}q$-orbit, see Definition 15. Thus, $\varepsilon\in\eta_{L}^{-1}(s)$ which implies that $s=\eta_{L}(\varepsilon)=i$. Hence, $i=s$ and consequently, $(is)^{n}=(ii)^{n}=i=(ii)^{n+1}=(is)^{n+1}$ which contradicts our assumption that $is$ is periodic.
We now continue with two cases, with the knowledge that $\varrho(\eta_{L}^{-1}(s))\neq\emptyset$ and $\varrho(\eta_{L}^{-1}(i))\neq\emptyset$:
Case 1, $|\varrho(\eta_{L}^{-1}(s))\mathrel{\cup}\varrho(\eta_{L}^{-1}(i))|>1$
Let $u\in\eta_{L}^{-1}(i)$ and $v\in\eta_{L}^{-1}(s)$ such that $\varrho(u)\neq\varrho(v)$. Since $i$ is idempotent, we have that $uuv,uv\in\eta_{L}^{-1}(is)$, as $iis=is$. We know that $is\in M_{L}$ is periodic, and thus proving $\varrho(uuv)\neq\varrho(uv)$ would contradict our assumption that $\eta_{L}$ is group primitive. Therefore, the rest of the proof of this case is dedicated to proving $\varrho(uuv)\neq\varrho(uv)$: Working towards a contradiction, assume that $uv=r^{i}$ and $uuv=r^{j}$ for some $i,j\in\mathbb{N}_{+}$ and for some primitive word $r\in\Sigma^{+}$. Notice that $(uv)^{m}=r^{2ij}=(uuv)^{n}$ where $m=2j$ and $n=2i$. This implies that $v(uv)^{m-1}=uv(uuv)^{n-1}$. Taking the prefix of length $|uv|$ from the left and right-hand side of the previous equality gives us $vu=uv$. Proposition 1.3.2 from [26] states that $vu=uv$ implies $\varrho(u)=\varrho(v)$, which is a contradiction.
Case 2, $|\varrho(\eta_{L}^{-1}(s))\mathrel{\cup}\varrho(\eta_{L}^{-1}(i))|=1$
Let $w\in\Sigma^{+}$ such that $\varrho(\eta_{L}^{-1}(s))=\{w\}=\varrho(\eta_{L}^{-1}(i))$. Since $\eta_{L}^{-1}(i)\subseteq w^{*}$ is regular, we know that $\eta_{L}^{-1}(i)\in\mathcal{B}q$. Thus, there exists $L^{\prime}\in\mathcal{B}q$ such that $\eta_{L}^{-1}(i)=L^{\prime}$ (which clearly implies $\eta_{L}^{-1}(i)\subseteq L^{\prime}$). Now, by the definition of a $\mathcal{B}q$-orbit, $\eta_{L}^{-1}(s)\mathrel{\cap}L^{\prime}\neq\emptyset$. Thus, $u\in\eta_{L}^{-1}(s)\mathrel{\cap}\eta_{L}^{-1}(i)$ for some $u\in\Sigma^{*}$. This implies that $\eta_{L}(u)=s=\eta_{L}(u)=i$. Consequently, $i=s$, and hence $(is)^{n}=(ii)^{n}=(ii)^{n+1}=(is)^{n+1}$ as $i$ is idempotent. This contradicts our assumption that $is\in M$ is periodic.
Immediately from Theorem 16, Lemma 17 and Lemma 18, we get the main result of this section:
Theorem 19.
If $L\subseteq\Sigma^{*}$ is group primitive, then $L\in\mathsf{SF}(\mathcal{R})$.
Proof:
Let $\eta_{L}\colon\Sigma^{*}\rightarrow M_{L}$ be a group primitive syntactic morphism of the regular language $L\subseteq\Sigma^{*}$. By Lemma 18, we know that all $\mathcal{B}q$-orbits with respect to $\mu$ are aperiodic monoids. Invoking Theorem 16, we know that $L\in\mathsf{SF}(\mathcal{B}q)$, and by Lemma 17 this implies $L\in\mathsf{SF}(\mathcal{R})$.
Summarising the results from Section V and this section, we have that $\mathsf{FC}$ can define all the languages in $\mathsf{SF}(\mathcal{R})$, and that if a language is a group primitive language, then it can be expressed in $\mathsf{SF}(\mathcal{R})$. An immediate corollary is that any group primitive language can be expressed in $\mathsf{FC}$.
VII On Loop-Step Cycles and $\mathsf{FC}$-Definability
In this section, we show that any regular language whose minimal DFA has a loop-step cycle is not expressible in $\mathsf{FC}$. In order to achieve inexpressibility results for $\mathsf{FC}$, we use a result from Lynch [24], which provides sufficient conditions for Duplicator to have a winning strategy for Ehrenfeucht-Fraïssé games over structures of the form $(\mathbb{N},+,A)$ where $A\subseteq\mathbb{N}$. Then, we shall generalize these inexpressibility results to prove that if an automaton has a loop-step cycle (Definition 7), then the language of that automaton is not definable in $\mathsf{FC}$.
VII-A Some Background on Ehrenfeucht-Fraïssé Games
The following definitions concerning Ehrenfeucht-Fraïssé games follow closely to the definitions given in Chapter 3 of Libkin [20]. Ehrenfeucht-Fraïssé-games are played by two players called Spoiler and Duplicator. The board consists of two relational structures $\mathcal{A}$ and $\mathcal{B}$ over the same signature. They play for a predetermined number $k\in\mathbb{N}$ of rounds, and each round is as follows:
  • Spoiler picks a structure $\mathcal{A}$ or $\mathcal{B}$, and then picks some element of their chosen structure.
  • Duplicator responds by picking some element in the structure that Spoiler did not choose.
In order to define winning conditions, we first define a partial isomorphism.
Definition 20 (Partial Isomorphism).
Let $\mathcal{A}$ and $\mathcal{B}$ be two relational structures over the same signature $\tau$, with the universes $A$ and $B$ respectively. Let $\vec{a}=(a_{1},\dots,a_{n})\in A^{n}$ and $\vec{b}=(b_{1},\dots,b_{n})\in B^{n}$ Then $(\vec{a},\vec{b})$ defines a partial isomorphism between $\mathcal{A}$ and $\mathcal{B}$ if:
  • For every $i,j\in[n]$, we have that $a_{i}=a_{j}$ iff $b_{i}=b_{j}$;
  • For every constant symbol $c\in\tau$ and every $i\in[n]$ we have that $a_{i}=c^{\mathcal{A}}$ iff $b_{i}=c^{\mathcal{B}}$.
  • For every relation symbol $R\in\tau$ with arity $m$, and every sequence $(i_{1},\dots,i_{m})\in[n]^{m}$, we have that $(a_{i_{1}},\dots,a_{i_{m}})\in R^{\mathcal{A}}$ iff $(b_{i_{1}},\dots,b_{i_{m}})\in R^{\mathcal{B}}$.
Let $\tau$ be a relational signature with constant symbols $c_{1},\dots,c_{r}$. Let $\mathcal{A}$ and $\mathcal{B}$ be two $\tau$-structures, and let $\mathcal{G}$ denote the $k$-round Ehrenfeucht-Fraïssé game over $\mathcal{A}$ and $\mathcal{B}$. Then, the resulting tuples of $\mathcal{G}$ are $\vec{a}=(a_{1},\dots,a_{k},c_{1}^{\mathcal{A}},\dots,c_{r}^{\mathcal{A}})$ and $\vec{b}=(b_{1},\dots,b_{k},c_{1}^{\mathcal{B}},\dots,c_{r}^{\mathcal{B}})$ where for all $i\in[k]$, we have that $a_{i}\in A$ is the element of $\mathcal{A}$ chosen in round $i$, and $b_{i}\in B$ is the element of $\mathcal{B}$ chosen in round $i$. Then, Duplicator wins $\mathcal{G}$ if and only if $(\vec{a},\vec{b})$ forms a partial isomorphism.
For a $k$-round game $\mathcal{G}$ over $\mathcal{A}$ and $\mathcal{B}$, we say that Duplicator has a winning strategy for $\mathcal{G}$ if and only if there is a way for Duplicator to play such that no matter what Spoiler chooses, Duplicator can win $\mathcal{G}$. Otherwise, Spoiler has a winning strategy for $\mathcal{G}$. If Duplicator has a winning strategy for $\mathcal{G}$, then we write $\mathcal{A}\equiv_{k}\mathcal{B}$. If $\mathcal{A}$ and $\mathcal{B}$ are $\mathsf{FC}$-structures, then $\mathcal{A}$ and $\mathcal{B}$ are uniquely determined (up to isomorphism) by some $w,v\in\Sigma^{*}$. Therefore, we can simply write $w\equiv_{k}v$.
Each first-order formula $\varphi$ has a so-called quantifier rank, denoted by the function $\mathsf{qr}$. Let $\varphi,\psi$ be first-order formulas over the same signature. We define $\mathsf{qr}$ recursively as follows:
  • If $\varphi\in\mathsf{FC}$ is an atomic formula, then $\mathsf{qr}(\varphi):=0$,
  • $\mathsf{qr}(\neg\varphi):=\mathsf{qr}(\varphi)$,
  • $\mathsf{qr}(\varphi\land\psi)=\mathsf{qr}(\varphi\lor\psi):=\mathsf{max}\left(\mathsf{qr}(\varphi),\mathsf{qr}(\psi)\right)$, and
  • $\mathsf{qr}(Qx\colon\varphi):=\mathsf{qr}(\varphi)+1$ for any $x\in X$ and $Q\in\{\forall,\exists\}$.
Let $\mathsf{FC}(k)$ denote the set of sentences $\varphi\in\mathsf{FC}$ such that $\mathsf{qr}(\varphi)\leqslant k$. Note if $w\not\equiv_{k}v$ for $w,v\in\Sigma^{*}$, then there exists $\varphi\in\mathsf{FC}(k)$ where $w\models\varphi$ and $v\not\models\varphi$; see Theorem 3.3 of [11].
It is known that winning strategies for Ehrenfeucht-Fraïssé games completely characterize the expressive power of first-order logic. We state this correspondence using a slight adaptation of Theorem 6.19 from Immerman [31].
Theorem 21.
Let $\mathcal{C}$ be a class of finite or infinite structures over a finite signature, and let $P\subseteq\mathcal{C}$. $P$ is not expressible in first-order logic if and only if for all $k\in\mathbb{N}$, there exists $\mathcal{A},\mathcal{B}\in\mathcal{C}$ such that:
  1. 1.
    $\mathcal{A}\in P$ and $\mathcal{B}\notin P$, and
  2. 2.
    $\mathcal{A}\equiv_{k}\mathcal{B}$.
As $\mathsf{FC}$-structures are simply a class of structures over the signature $\tau_{\Sigma}$, it follows that Theorem 21 immediately holds for $\mathsf{FC}$ (see [11] for more details on Ehrenfeucht-Fraïssé games over $\mathsf{FC}$-structures).
VII-B A Regular Language Not Expressible in $\mathsf{FC}$
For our next step, we utilize Lynch’s Theorem [24], which gives sufficient conditions for Duplicator to have a winning strategy over certain structures with addition. We then connect this with $\mathsf{FC}$ by encoding these structures as words which allows us to adapt the winning strategies given by Lynch’s Theorem to inexpressibility results for $\mathsf{FC}$.
By $(\mathbb{N},+,A)$, we denote the structure with the universe $\mathbb{N}$, an addition relation $z=x+y$, and a finite set $A\subseteq\mathbb{N}$ [^4] We note that Lynch [24] did not require $A$ to be finite.
Definition 22 (Lynch [24]).
Let $d$, $f$, and $g$ be functions from $\mathbb{N}$ to $\mathbb{N}$ defined as follows:
  • $d(0)=5$, and $d(i+1)=(2^{i+3}+1)d(i)$,
  • $f(0)=1$, and $f(i+1)=2f(i)^{4}$, and
  • $g(0)=0$, and $g(i+1)=2f(i)^{2}g(i)+f(i)!$.
The exact definitions of $d$, $f$, and $g$ are not important for this work. We only give their definitions for the following:
Definition 23 (Adapted from Lynch [24]).
Let $k\in\mathbb{N}$ and let $(p_{i})_{i\in\mathbb{N}}$ be any sequence in $\mathbb{N}$ such that $p_{0}=0$ and
$p_{i+1}\geqslant 2^{k+3}f(k)^{3}p_{i}+2f(k)^{2}g(k)$ (1)
for $i\in\mathbb{N}$, and $p_{i}\equiv p_{j}\pmod{f(k)!}$ for $i,j\in\mathbb{N}_{+}$. Then, any set of the form $P_{k}=\{p_{i}\mid i\geqslant 1\}$ is a $k$-Lynchian set.
We are now ready to state Lynch’s Theorem:
Theorem 24 (Lynch’s Theorem [24]).
Let $k\in\mathbb{N}$ and let $P_{k}\subseteq\mathbb{N}$ be a $k$-Lynchian set. For any finite sets $A,B\subseteq P_{k}$ where $d(k)<|A|,|B|$ we have that $(\mathbb{N},+,A)\equiv_{k}(\mathbb{N},+,B)$.
Informally, Lynch’s Theorem states that as long as $A$ and $B$ are big enough, and only contain elements from a $k$-Lynchian set, we have that $(\mathbb{N},+,A)\equiv_{k}(\mathbb{N},+,B)$. As an example, for any $k\in\mathbb{N}$ we can pick $A$ and $B$ such that $|A|$ is even, $|B|$ is odd, and $(\mathbb{N},+,A)\equiv_{k}(\mathbb{N},+,B)$.
Our next focus is on adapting Ehrenfeucht-Fraïssé game strategies over structures with addition to $\mathsf{FC}$-structures. Thus gaining new insights into the expressive power of $\mathsf{FC}$. Namely, we shall prove that there exist some regular languages that cannot be expressed by $\mathsf{FC}$.
Definition 25.
For every finite set $A\subset\mathbb{N}$, the canonical encoding of $A=\{a_{1}<a_{2}<\dots<a_{n}\}$ as a word $\bar{w}_{A}\in\{\mathtt{0},\mathtt{1}\}^{*}$ is $\bar{w}_{A}:=\mathtt{1}\cdot\mathtt{0}^{a_{1}}\cdot\mathtt{1}\cdot\mathtt{0}^{a_{2}}\cdot\mathtt{1}\cdots\mathtt{1}\cdot\mathtt{0}^{a_{n}}\cdot\mathtt{1}$.
Let $\mathsf{FO}[+,S]$ denote first-order logic over the class of structures of the form $(\mathbb{N},+,S)$ where $S\subseteq\mathbb{N}$. Using the canonical encoding of a set as a word, we can translate formulas over $\mathsf{FC}$-structures to $\mathsf{FO}[+,S]$.
Lemma 26.
There exists a function $r\colon\mathbb{N}\to\mathbb{N}$ such that for every $k\in\mathbb{N}$ and all finite sets $A,B\subset\mathbb{N}$ the following holds: If $(\mathbb{N},+,A)\equiv_{r(k)}(\mathbb{N},+,B)$ then $\bar{w}_{A}\equiv_{k}\bar{w}_{B}$.
Proof:
Consider an arbitrary $k\in\mathbb{N}$ and arbitrary finite sets $A,B\subset\mathbb{N}$. Let $\mathcal{A}:=(\mathbb{N},+,A)$ and $\mathcal{B}:=(\mathbb{N},+,B)$.
To prove the result, we prove the contraposition, i.e., we show that if $\bar{w}_{A}\not\equiv_{k}\bar{w}_{B}$, then $\mathcal{A}\not\equiv_{r(k)}\mathcal{B}$, for a suitably chosen number $r(k)$ that does not depend on the sets $A,B$.
Assume that $\bar{w}_{A}\not\equiv_{k}\bar{w}_{B}$, i.e., there exists an $\mathsf{FC}$-sentence $\varphi\in\mathsf{FC}(k)$ such that $\bar{w}_{A}\models\varphi$ and $\bar{w}_{B}\not\models\varphi$. In the following, we construct a first-order reduction that translates $\varphi$ into a sentence $\psi\in\mathsf{FO}[+,S]$ such that $\mathcal{A}\models\psi$ and ⊧̸$\mathcal{B}\not\models\psi$. That is, using formulas in $\mathsf{FO}[+,S]$, we simulate all the behaviour of an $\mathsf{FC}$-formula over words that canonically encode a set (as in Definition 25). See Chapter 3 of Immerman [31] for more details on first-order reductions. Finally, we will be done by letting $r(k):=\mathsf{qr}(\psi)$ and by noting that the quantifier rank of $\psi$ only depends on $k$ and the first-order reduction, but is independent of the concrete formula $\varphi$.
We extend $\{+,S\}$-structures with a new constant $\bot$ which does not appear in the addition relation; clearly this does not change the expressive power of $\mathsf{FO}[+,S]$. Let $\bar{w}_{A}\in\{\mathtt{0},\mathtt{1}\}^{*}$ be the canonical encoding of a non-empty and finite set $A=\{x_{1}<x_{2}<\dots<x_{|A|}\}\subset\mathbb{N}$. For every factor $u\sqsubseteq\bar{w}_{A}$, we associate a tuple $t_{u}=(a_{1},a_{2},a_{3},a_{4})\in(\mathbb{N}\mathrel{\cup}\{\bot\})^{4}$:
  • If $|u|_{\mathtt{1}}\geqslant 2$, then $u=\mathtt{0}^{m}\cdot\mathtt{1}\cdot\mathtt{0}^{x_{l}}\cdot\mathtt{1}\cdots\mathtt{1}\cdot\mathtt{0}^{x_{r}}\cdot\mathtt{1}\cdot\mathtt{0}^{n}$, where $1\leqslant l\leqslant r\leqslant|A|$. Thus, we let $t_{u}:=(m,x_{l},x_{r},n)$.
  • If $|u|_{\mathtt{1}}=1$, then $u=\mathtt{0}^{m}\mathtt{1}\mathtt{0}^{n}$. Thus, $t_{u}:=(m,\bot,\bot,n)$.
  • If $|u|_{\mathtt{0}}=0$, then $u=\mathtt{0}^{m}$. Thus, $t_{u}:=(m,\bot,\bot,\bot)$.
Universe
Let $\psi_{\mathsf{univ}}(x_{1},x_{2},x_{3},x_{4})$ be a formula in $\mathsf{FO}[+,S]$ such that for any $\mathcal{A}$, we have
$\psi_{\mathsf{univ}}(\mathcal{A})=\{(a_{1},a_{2},a_{3},a_{4})\in(\mathbb{N}\mathrel{\cup}\{\bot\})^{4}\mid\\ u\sqsubseteq\bar{w}_{A}\text{ and }(a_{1},a_{2},a_{3},a_{4})\text{ is the tuple for }u\}.$
It is easy to see that $\psi_{\mathsf{univ}}(x_{1},x_{2},x_{3},x_{4})$ can be written in $\mathsf{FO}[+,S]$.
Constants
Since the constant symbol $\mathtt{0}$ is represented by the tuple $(1,\bot,\bot,\bot)$, let
$\psi_{\mathtt{0}}(x_{1},x_{2},x_{3},x_{4}):=(x_{1}\mathbin{\dot{=}}1)\land\bigwedge_{2\leqslant i\leqslant 4}(x_{i}\mathbin{\dot{=}}\bot).$
Since the constant symbol $\mathtt{1}$ is represented by $(0,\bot,\bot,0)$, let
$\psi_{\mathtt{1}}(x_{1},x_{2},x_{3},x_{4}):=(x_{1}\mathbin{\dot{=}}0)\land(x_{4}\mathbin{\dot{=}}0)\land(x_{3}\mathbin{\dot{=}}\bot)\land(x_{4}\mathbin{\dot{=}}\bot).$
Concatenation
To simulate concatenation, we define the ternary relation $\oplus$ over elements of $\psi_{\mathsf{univ}}(\mathcal{A})$. We wish to have that $t_{w}=t_{u}\oplus t_{v}$ if and only if $w=uv$. To realize this behaviour, we look at various cases. However, we shall only look at one case in detail, as the others follow analogously.
  • Case 1, $|u|_{\mathtt{1}},|v|_{\mathtt{1}}\geqslant 2$. For this case, we have that $u=\mathtt{0}^{m}\mathtt{1}\mathtt{0}^{x_{l}}\mathtt{1}\cdots\mathtt{1}\mathtt{0}^{x_{r}}\mathtt{1}\mathtt{0}^{n}$ and $v=\mathtt{0}^{m^{\prime}}\mathtt{1}\mathtt{0}^{x_{l^{\prime}}}\mathtt{1}\cdots\mathtt{1}\mathtt{0}^{x_{r^{\prime}}}\mathtt{1}\mathtt{0}^{n^{\prime}}$. Then, $u\cdot v=\mathtt{0}^{m}\mathtt{1}\mathtt{0}^{x_{l}}\mathtt{1}\cdots\mathtt{0}^{x_{r^{\prime}}}\mathtt{1}\mathtt{0}^{n^{\prime}}$ is a factor of $\bar{w}_{A}$, if $n+m^{\prime}=x_{r+1}=x_{l^{\prime}-1}$. Thus
    $\underbrace{(m,x_{l},x_{r^{\prime}},n^{\prime})}_{t_{uv}}=\underbrace{(m,x_{l},x_{r},n)}_{t_{u}}\oplus\underbrace{(m^{\prime},x_{l^{\prime}},x_{r^{\prime}},n^{\prime})}_{t_{v}}$
    whenever $n+m^{\prime}=x_{r+1}=x_{l^{\prime}-1}$. It is clear that one could write an $\mathsf{FO}[+,S]$-formula to realize this behaviour: Consider a formula with arity $12$, which first ensures the first four, the middle four, and the last four components satisfy $\psi_{\mathsf{univ}}(\mathcal{A})$. The formula also ensures that none of the componts are $\bot$, and the required arithmetic and equalities holds.
  • Case 2, $|u|_{\mathtt{1}}\geqslant 2$ and $|v|_{\mathtt{1}}=1$. If $n+m^{\prime}=x_{r+1}$, then
    $(m,x_{l},x_{r+1},n^{\prime})=(m,x_{l},x_{r},n)\oplus(m^{\prime},\bot,\bot,n^{\prime}).$
  • Case 3, $|u|_{\mathtt{1}}=1$ and $|v|_{\mathtt{1}}\geqslant 2$. Symmetric to Case 2.
  • Case 4, $|u|_{\mathtt{1}}\geqslant 2$ and $|v|_{\mathtt{1}}=0$. If $n+m^{\prime}\leqslant x_{r+1}$, then
    $(m,x_{l},x_{r},n+m^{\prime})=(m,x_{l},x_{r},n)\oplus(m^{\prime},\bot,\bot,\bot).$
  • Case 5, $|u|_{\mathtt{1}}=0$ and $|v|_{\mathtt{1}}\geqslant 2$. Symmetric to Case 4.
  • Case 6, $|u|_{\mathtt{1}}=1$ and $|v|_{\mathtt{1}}=1$. If $n+m^{\prime}=x_{i}\in A$ where $m\leqslant x_{i-1}$ and $n^{\prime}\leqslant x_{i+1}$, then let
    $(m,n+m^{\prime},n+m^{\prime},n^{\prime})=(m,\bot,\bot,n)\oplus(m^{\prime},\bot,\bot,n^{\prime}).$
  • Case 7, $|u|_{\mathtt{1}}=1$ and $|v|_{\mathtt{1}}=0$. If $n+m^{\prime}\leqslant\mathsf{max}(A)$, then
    $(m,\bot,\bot,n+m^{\prime})=(m,\bot,\bot,n)\oplus(m^{\prime},\bot,\bot,\bot).$
  • Case 8, $|u|_{\mathtt{1}}=0$ and $|v|_{\mathtt{1}}=1$. Symmetric to Case 7.
  • Case 9, $|u|_{\mathtt{1}}=0$ and $|v|_{\mathtt{1}}=0$. If $n+m^{\prime}\leqslant\mathsf{max}(A)$, then
    $(m+m^{\prime},\bot,\bot,\bot)=(m,\bot,\bot,\bot)\oplus(m^{\prime},\bot,\bot,\bot).$
It is a straightforward exercise to define a $\mathsf{FO}[+,S]$-formula for each of the above cases. Combining the formulas for each case with disjunction results in a formula that has the correct behaviour for $\oplus$. That is, it is easily observed that $t_{uv}=t_{u}\oplus t_{v}$ does indeed hold with the above definition of $\oplus$.
One can now rewrite $\varphi\in\mathsf{FC}(k)$ that separates $\bar{w}_{A}$ and $\bar{w}_{B}$ in a straightforward manner, using the above defined concepts, into a formula $\psi\in\mathsf{FO}[+,S]$ such that $\psi$ separates $\mathcal{A}$ and $\mathcal{B}$. More formally, we shall define a mapping $\mathcal{T}\colon\mathsf{FC}\rightarrow\mathsf{FO}[+,S]$. First, let $\ell\colon X\mathrel{\cup}\{\mathtt{0},\mathtt{1}\}\rightarrow(\Gamma\mathrel{\cup}\{\mathtt{0},\mathtt{1},\bot\})^{4}$ be a function that maps each $\mathsf{FC}$ variable $x\in X$ to a unique tuple of variables $\vec{x}\in\Gamma^{4}$ (where $\Gamma$ is a countably infinite set of variables that is disjoint from $X$), maps $\mathtt{0}$ to $(1,\bot,\bot,\bot)$ and maps $\mathtt{1}$ to $(0,\bot,\bot,0)$. Note that as $\ell(\mathtt{0})$ and $\ell(\mathtt{1})$ can easily be expressed with an $\mathsf{FO}[+,S]$-formula, we can treat them as new constant symbols. Likewise, we treat $\oplus$ as a relational symbol as it can also be expressed in $\mathsf{FO}[+,S]$. For $\varphi,\varphi^{\prime}\in\mathsf{FC}$, let
  • $\mathcal{T}(Qx\colon\varphi):=Q\vec{x}\colon(\psi_{\mathsf{univ}}(\vec{x})\land\mathcal{T}(\varphi))$ for $Q\in\{\forall,\exists\}$,
  • $\mathcal{T}(\varphi\land\varphi^{\prime}):=\mathcal{T}(\varphi)\land\mathcal{T}(\varphi^{\prime})$,
  • $\mathcal{T}(\varphi\lor\varphi^{\prime}):=\mathcal{T}(\varphi)\lor\mathcal{T}(\varphi^{\prime})$,
  • $\mathcal{T}(\neg\varphi):=\neg\mathcal{T}(\varphi)$,
  • $\mathcal{T}(x\mathbin{\dot{=}}y\cdot z):=\bigl{(}\ell(x)\mathbin{\dot{=}}\ell(y)\oplus\ell(z)\bigr{)}$.
Here, if $\vec{x}=(y_{1},y_{2},y_{3},y_{4})$ and $Q\in\{\exists,\forall\}$, we use $Q\vec{x}$ as shorthand for $Qy_{1},y_{2},y_{3},y_{4}$.
Then, for any $\varphi\in\mathsf{FC}$ such that $\bar{w}_{A}\models\varphi$ and $\bar{w}_{B}\not\models\varphi$, we have that $\mathcal{A}\models\mathcal{T}(\varphi)$ and ⊧̸$\mathcal{B}\not\models\mathcal{T}(\varphi)$. Consequently, there exists some $\mathsf{FO}[+,S]$-formula $\psi$ that separates $\mathcal{A}$ from $\mathcal{B}$, which means that $\mathcal{A}\not\equiv_{r(k)}\mathcal{B}$ where $r(k)=\mathsf{qr}(\psi)$.
By combining Theorem 24 and Lemma 26, we obtain that there are regular languages that cannot be expressed in $\mathsf{FC}$.
Theorem 27.
$L:=\{w\in\{\mathtt{0},\mathtt{1}\}^{*}\mid|w|_{\mathtt{1}}\text{ is even}\}$ is a regular language that is not $\mathsf{FC}$-definable.
Proof:
By Theorem 21 it suffices to construct, for every $k\in\mathbb{N}$, two words $u_{k}\in L$ and $v_{k}\not\in L$ such that $u_{k}\equiv_{k}v_{k}$.
Let us fix an arbitrary $k\in\mathbb{N}$, and let $k^{\prime}:=r(k)$ be the number provided by Lemma 26. Let $P_{k^{\prime}}\subseteq\mathbb{N}$ be a $k^{\prime}$-Lynchian set. From Theorem 24 we know that for finite sets $A,B\subseteq P_{k^{\prime}}$ where $d(k^{\prime})<|A|,|B|$ we have $(\mathbb{N},+,A)\equiv_{k^{\prime}}(\mathbb{N},+,B)$. Thus, the only restriction on $|A|$ and $|B|$ is that they are large enough. Here, we choose finite sets $A,B\subseteq P_{k^{\prime}}$ where where $|A|$ is even and $|B|$ is odd and where $d(k^{\prime})<|A|,|B|$. It follows from Theorem 24 that $(\mathbb{N},+,A)\equiv_{k^{\prime}}(\mathbb{N},+,B)$. As $k^{\prime}=r(k)$, by Lemma 26 we obtain: $\bar{w}_{A}\equiv_{k}\bar{w}_{B}$, where $\bar{w}_{A},\bar{w}_{B}\in\{\mathtt{0},\mathtt{1}\}^{*}$ are the canonical encodings of $A$ and $B$ respectively.
Notice since $|A|$ is even and $|B|$ is odd, $|\bar{w}_{A}|_{\mathtt{1}}$ is odd and $|\bar{w}_{B}|_{\mathtt{1}}$ is even. Thus, $\bar{w}_{B}\in L$, $\bar{w}_{A}\not\in L$ and $\bar{w}_{B}\equiv_{k}\bar{w}_{A}$.
The technique used in the above proof (for different choices of the sets $A,B$) actually give a range of further regular languages that are not $\mathsf{FC}$-definable. In particular, the connection between addition and concatenation provided by Lemma 26 is a key step that we will also use in the following subsections in order to prove that the language of any minimal DFA that has a loop-step cycle is not $\mathsf{FC}$-definable.
VII-C Generalizing With Morphisms
Our next step is to generalize Lemma 26 using morphisms of the form $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$. Informally, our goal is to show that $\bar{w}_{A}\equiv_{k}\bar{w}_{B}$ implies $h(\bar{w}_{A})\equiv_{k^{\prime}}h(\bar{w}_{B})$, for a specific type of morphism $h$.
We start this section with a folklore result, see Section 2.2, Chapter 6 of [32] for more details of Lemma 28:
Lemma 28 (Folklore).
A word $w\in\Sigma^{+}$ is an internal factor of $ww$ if and only if $w$ is imprimitive.
For some of the subsequent results, we require concepts from coding theory. Most of the subsequent definitions are reformulations of definitions from Chapter 6 of [33].
Definition 29.
A set $\Lambda\subseteq\Sigma^{*}$ is a code if
$x_{1}\cdot x_{2}\cdots x_{n}=y_{1}\cdot y_{2}\cdots y_{m}$
with $x_{1},\dots,x_{n},y_{1},\dots,y_{m}\in\Lambda$ implies that $n=m$ and $x_{i}=y_{i}$ for $i\in[n]$.
A slight rephrasing of Definition 29 gives us the following:
Lemma 30 (Proposition 6.1.3 [33]).
A set $\Lambda\subseteq\Sigma^{+}$ is a code if and only if any morphism $h\colon A^{*}\rightarrow\Sigma^{*}$ induced by a bijection from $A$ onto $\Lambda$ is injective.
A set $\{u,v\}$ with $u,v\in\Sigma^{*}$ is bifix if $u$ is not a prefix or suffix of $v$ and vice versa, see Chapter 6 of [34]. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a non-erasing morphism where $\varrho(h(\mathtt{0}))\neq\varrho(h(\mathtt{1}))$. We say that $h$ is a bifix morphism if the set $\{h(\mathtt{0}),h(\mathtt{1})\}$ is bifix.
Lemma 31.
Let $|\Sigma|\geqslant 2$. If $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ is a bifix morphism, then $\{h(\mathtt{0}),h(\mathtt{1})\}$ is a code.
Proof:
This follows from Proposition 2.1.9 of [34] which states that any non-empty set of words $\Lambda\neq\{\varepsilon\}$ such that $\Lambda\mathrel{\cap}\Lambda\Sigma^{+}=\emptyset$ is a code.
Immediately from Lemma 31 and Lemma 30 we know that a bifix morphism $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ is injective.
Definition 32.
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a morphism and let $w\in\{\mathtt{0},\mathtt{1}\}^{*}$. Then, for any $u\sqsubseteq h(w)$, we say that $(x,y,z)\in\Sigma^{*}\times\{\mathtt{0},\mathtt{1}\}^{*}\times\Sigma^{*}$ is a core factorization of $u$ with respect to $w$ and $h$ if $u=x\cdot h(y)\cdot z$ and:
  • $x\sqsubset_{\mathsf{suff}}h(c)$ for some $c\in\{\mathtt{0},\mathtt{1}\}$,
  • $y\sqsubseteq w$, and
  • $z\sqsubset_{\mathsf{pref}}h(c^{\prime})$ for some $c^{\prime}\in\{\mathtt{0},\mathtt{1}\}$.
If $(x,y,z)$ is a core factorization of $u$ with respect to $w$ and $h$, then $y\sqsubseteq w$ is a pre-image core of $u$ with respect to $w$ and $h$.
When $w\in\Sigma^{*}$ and $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ are clear from context, we simply say $(x,y,z)$ is a core factorization of $u$, and $y$ is a pre-image core of $u$.
Recall Definition 25 of a canonical encoding of a finite set $A\subset\mathbb{N}$ as a word $\bar{w}_{A}\in\{\mathtt{0},\mathtt{1}\}^{*}$. As a quick example, the set $A=\{1,2,4\}$ is encoded as $\bar{w}_{A}=\mathtt{1}\cdot\mathtt{0}\cdot\mathtt{1}\cdot\mathtt{0}^{2}\cdot\mathtt{1}\cdot\mathtt{0}^{4}\cdot\mathtt{1}$.
Lemma 33.
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a bifix morphism where $h(\mathtt{0})$ is primitive and $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$. Let $\bar{w}_{A}\in\{\mathtt{0},\mathtt{1}\}^{*}$ be the canonical encoding of a set $A\subset\mathbb{N}$ where $\mathsf{min}(A)\geqslant 2|h(\mathtt{1})|$. Then, for any $u\sqsubseteq h(\bar{w}_{A})$ where $|u|>10\cdot\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)$, the pre-image core of $u$ is unique.
Proof:
We assume there exist two distinct core factorizations for $u$ with respect to $\bar{w}_{A}$ and $h$, and show a contradiction. Namely, either that $h$ is not bifix, or $h(\mathtt{0})$ is not primitive (utilizing Lemma 28 and Lemma 31).
For our next step, we utilise a basic lemma from [11]. Informally, this result states that if Spoiler picks a short factor (with respect to the number of remaining rounds), then Duplicator must respond with the identical factor (or lose).
Lemma 34 (Thompson and Freydenberger [11]).
Let $\mathcal{A}_{w}$ and $\mathcal{B}_{v}$ be $\tau_{\Sigma}$-structures that represent $w\in\Sigma^{*}$ and $v\in\Sigma^{*}$, where $\mathcal{A}_{w}\equiv_{k}\mathcal{B}_{v}$. Let $\vec{a}=(a_{1},a_{2},\dots,a_{k+|\Sigma|+1})$ and $\vec{b}=(b_{1},b_{2},\dots,b_{k+|\Sigma|+1})$ be the tuple resulting from a $k$-round game over $\mathcal{A}_{w}$ and $\mathcal{B}_{v}$ where Duplicator plays their winning strategy. If $|a_{r}|<k-r+1$ or $|b_{r}|<k-r+1$ for some $r\in[k]$, then $b_{r}=a_{r}$.
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a bifix morphism where $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$ and where $h(\mathtt{0})$ is primitive. Let $\bar{w}_{A}\in\{\mathtt{0},\mathtt{1}\}^{*}$ be the canonical encoding of a set $A\subset\mathbb{N}$ where the $\mathsf{min}(A)>2\cdot|h(\mathtt{1})|$. Then, from Lemma 33, we know that the core-factorization of $u\sqsubseteq h(\bar{w}_{A})$ is unique. This gives us a way to associate every long enough factor of $h(\bar{w}_{A})$ with some factor of $\bar{w}_{A}$, and hence gives a tool for generalizing Duplicator’s strategy over words of the form $\bar{w}_{A}$.
Lemma 35.
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a bifix morphism where $h(\mathtt{0})$ is primitive, and $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$. If $\bar{w}_{A}\equiv_{k+20}\bar{w}_{B}$ for sets $A,B\subseteq\mathbb{N}$ where $\mathsf{min}(A\mathrel{\cup}B)>2\cdot\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)+10$, then $h(\bar{w}_{A})\equiv_{k}h(\bar{w}_{B})$.
Proof:
To prove this result, we use Duplicator’s winning strategy for the Ehrenfeucht-Fraïssé game $\mathcal{G}^{\prime}$ over $\bar{w}_{A}$ and $\bar{w}_{B}$ as a “subroutine” for a Duplicator’s strategy for the game $\mathcal{G}$ over $h(\bar{w}_{A})$ and $h(\bar{w}_{B})$. Let $\mathsf{m}$ denote $\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)$. Assuming Spoiler chooses $a_{i}\sqsubseteq h(\bar{w}_{A})$ in round $i$, Duplicator’s response is as follows:
  1. 1.
    If $|a_{i}|>10\mathsf{m}$, then let $(u_{i},c_{i},u_{i}^{\prime})$ be the unique core factorization of $a_{i}$. Let Spoiler choose $c_{i}$ in round $i$ of $\mathcal{G}^{\prime}$ and let $d_{i}$ be Duplicator’s response in $\mathcal{G}^{\prime}$. Then, Duplicator’s response in $\mathcal{G}$ is $b_{i}:=u_{i}\cdot h(d_{i})\cdot u_{i}^{\prime}$.
  2. 2.
    If $|a_{i}|\leqslant 10\mathsf{m}$ then let Duplicator respond with $b_{i}:=a_{i}$. We assume that Spoiler chooses $\varepsilon$ in $\mathcal{G}^{\prime}$ for round $i$; the structure Spoiler chooses here is not important.
The main difficulty in this proof is showing that this strategy is indeed a winning strategy.
Lemma 35 is rather technical and specific to our purposes. It gives us a tool to generalize Duplicator’s strategy from Lemma 26 using specific morphisms $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$. We shall see in the subsequent section that this is a key step in proving the language of a minimal DFA is not $\mathsf{FC}$-definable if it has a loop-step cycle.
VII-D Loop-Step Cycles and Non-Definability in $\mathsf{FC}$
This section gives the actual proof that if a minimal DFA $\mathcal{M}$ has a loop-step cycle, then $\mathcal{L}(\mathcal{M})\notin\mathcal{L}(\mathsf{FC})$.
We rely on the following result:
Lemma 36 (Thompson and Freydenberger [11]).
Let $w_{1},w_{2},v_{1},v_{2}\in\Sigma^{*}$ where $\mathsf{facts}(w_{1})\mathrel{\cap}\mathsf{facts}(w_{2})$ is equal to $\mathsf{facts}(v_{1})\mathrel{\cap}\mathsf{facts}(v_{2})$, and let $r$ be the length of the longest word in $\mathsf{facts}(w_{1})\mathrel{\cap}\mathsf{facts}(w_{2})$. If $w_{1}\equiv_{k+r+2}v_{1}$ and $w_{2}\equiv_{k+r+2}v_{2}$ for some $k\in\mathbb{N}_{+}$, then $w_{1}\cdot w_{2}\equiv_{k}v_{1}\cdot v_{2}$.
We are now ready for the main result of this section.
Theorem 37.
Let $\mathcal{M}:=(Q,\Sigma,\delta,q_{0},F)$ be a minimal DFA. If $\mathcal{M}$ has a loop-step cycle, then $\mathcal{L}(\mathcal{M})\notin\mathcal{L}(\mathsf{FC})$.
Proof:
Let $\mathcal{M}:=(Q,\Sigma,\delta,q_{0},F)$ be a minimal DFA that has a loop-step cycle, i.e., there exist $n\geqslant 2$ pairwise distinct states $p_{0},p_{1},\dots,p_{n-1}$ and words $w,v\in\Sigma^{+}$ where $\varrho(w)\neq\varrho(v)$ and where:
  • $\delta^{*}(p_{i},w)=p_{i}$ for all $i\in\{0,\dots,n-1\}$,
  • $\delta^{*}(p_{i},v)=p_{i+1\pmod{n}}$.
Note that since $\mathcal{M}$ is minimal, all states must be reachable, and thus there exists some word $p\in\Sigma^{*}$ such that $\delta^{*}(q_{0},p)=p_{0}$. Furthermore, there must exist some word $s\in\Sigma^{*}$ such that $\delta^{*}(p_{j},s)\in F$ for some $j\in\{0,\dots,n-1\}$ and $\delta^{*}(p_{\ell},s)\notin F$ for some $\ell\in\{0,\dots,n-1\}$ (the reason is that otherwise, we could combine states to get a smaller automaton that accepts the same language, contradicting the minimality of $\mathcal{M}$; for details see Section 4.3 in Chapter 2 of [32]).
We choose such words $p$ and $s$ and indices $0\leqslant j,\ell<n$ and keep them fixed throughout the remainder of this proof. I.e., we have: $\delta^{*}(q_{0},p)=p_{0}$ and $\delta^{*}(p_{j},s)\in F$ and $\delta^{*}(p_{\ell},s)\not\in F$.
Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ where:
  • $h(\mathtt{0}):=w^{n|v|}\cdot v^{n|w|+n+1}$, and
  • $h(\mathtt{1}):=w^{2n|v|}\cdot v^{n+1}$.
Observe that $|h(\mathtt{0})|=|h(\mathtt{1})|$, because
  • $|h(\mathtt{0})|=2n|v||w|+n|v|+|v|$, and
  • $|h(\mathtt{1})|=2n|v||w|+n|v|+|v|$.
Notice that $\delta^{*}(p_{i},h(\mathtt{0}))=\delta^{*}(p_{i},h(\mathtt{1}))=p_{i+1\pmod{n}}$ for any $i\in\{0,\dots,n{-}1\}$.
Theorem 16 of [35] states that for two primitive words $x,y\in\Sigma^{+}$ where $x\neq y$, the word $x^{n}y^{m}$ is primitive for any $n,m\geqslant 2$. From $\varrho(w)\neq\varrho(v)$ we thus obtain that $h(\mathtt{0})$ and $h(\mathtt{1})$ are two distinct primitive words. As they are of equal length, $\{h(\mathtt{0}),h(\mathtt{1})\}$ is bifix, and $h$ is a bifix morphism. Furthermore, $h(\mathtt{0})$ is primitive and $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$. Therefore, we can apply Lemma 35 for any $k\in\mathbb{N}$ and any choice of finite sets $A,B\subset\mathbb{N}$ whose minimal elements are sufficiently large.
Recall that $\delta^{*}(p_{i},h(\mathtt{1}))=p_{i+1\pmod{n}}$ for every $i$ where $0\leqslant i<n$. Furthermore, note that $\delta^{*}(p_{i},h(\mathtt{0})^{x})=p_{i}$ for every $x\in\mathbb{N}$ with $x\equiv 0\pmod{n}$.
For any finite set $S\subset\mathbb{N}$ such that $x\equiv 0\pmod{n}$ holds for all $x\in S$, consider the word
$\alpha_{S}:=p\cdot h(\bar{w}_{S})\cdot s.$
Note that the word $\bar{w}_{S}$ contains $|S|{+}1$ occurrences of the letter $\mathtt{1}$, and between any two occurrences of the letter $\mathtt{1}$, there are $x$ occurrences of the letter $\mathtt{0}$, for some $x\in S$. Thus, for every $r\in\{0,\ldots,n{-}1\}$ the following is true:
If $|S|{+}1\equiv r\pmod{n}$, then $\delta^{*}(p_{0},h(\bar{w}_{S}))=p_{r}$.
Due to our particular choice of the words $s,p$ and the indices $j,\ell$, we therefore obtain:
  • If $|S|{+}1\equiv j\pmod{n}$, then $\delta^{*}(q_{0},\alpha_{S})\in F$, and therefore $\alpha_{S}\in\mathcal{L}(\mathcal{M})$.
  • If $|S|{+}1\equiv\ell\pmod{n}$, then $\delta^{*}(q_{0},\alpha_{S})\not\in F$, and therefore $\alpha_{S}\not\in\mathcal{L}(\mathcal{M})$.
Recall that our ultimate goal is to show that $L:=\mathcal{L}(\mathcal{M})$ is not $\mathsf{FC}$-definable. Due to Theorem 21 it suffices to construct for every $k\in\mathbb{N}$ two words $u\in L$ and $u^{\prime}\not\in L$ such that $u\equiv_{k}u^{\prime}$.
Let us fix an arbitrary $k\in\mathbb{N}$, choose a suitable number $m$, and let $k^{\prime}:=r(k{+}m{+}20)$ be the number provided by Lemma 26. Let $P_{k^{\prime}}\subseteq\mathbb{N}$ be a $k^{\prime}$-Lynchian set, which satisfies the additional condition that $x\equiv 0\pmod{n}$ for all $x\in P_{k^{\prime}}$, note that according to Definition 23, this condition can easily be satisfied.
From Theorem 24 we know that $(\mathbb{N},+,A)\equiv_{k^{\prime}}(\mathbb{N},+,B)$ for all finite sets $A,B\subseteq P_{k^{\prime}}$ where $d(k^{\prime})<|A|,|B|$. As $k^{\prime}$ is $r(k{+}m{+}20)$, by Lemma 26 we obtain: $\bar{w}_{A}\equiv_{k+m+20}\bar{w}_{B}$. Provided that the minimal elements in $A$ and $B$ are sufficiently large, Lemma 35 yields that $h(\bar{w}_{A})\equiv_{k+m}h(\bar{w}_{B})$. Using Lemma 36, and having chosen $m$ sufficiently large, we obtain that $p\cdot h(\bar{w}_{A})\cdot s\equiv_{k}p\cdot h(\bar{w}_{B})\cdot s$, i.e., we have $\alpha_{A}\equiv_{k}\alpha_{B}$.
Finally, we complete this proof by choosing two finite sets $A,B\subseteq P_{k^{\prime}}$ where $d(k^{\prime})<|A|,|B|$, and the minimal elements in $A$ and $B$ are large enough for Lemma 35, and where we have $|A|{+}1\equiv j\pmod{n}$ and $|B|{+}1\equiv\ell\pmod{n}$. Then, for the words $u:=\alpha_{A}$ and $u^{\prime}:=\alpha_{B}$ we have: $u\in L$, $u^{\prime}\not\in L$, and (provided that we have chosen $m$ large enough for applying Lemma 36) $u\equiv_{k}u^{\prime}$. This proves that $L$ is not definable in $\mathsf{FC}$, and it completes the proof of Theorem 37.
[…]
VIII Conclusions
In this paper, we have provided a decidable characterization of the $\mathsf{FC}$-definable regular languages in terms of (generalized) regular expressions, automata, and algebra (Theorem 10). Moreover, we have shown that this characterization is decidable, and is in fact $\mathsf{PSPACE}$-complete for minimal DFAs (Theorem 11). A promising next step would be to study the expressive power of fragments of $\mathsf{FC}$. The existential-positive fragment is particularly interesting due to its tight connection with core spanners [1].
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable feedback and suggestions. This work was supported by the EPSRC grant EP/T033762/1.
Data Availability
No data was analysed, captured or generated for this work.
Appendix A Proof of Theorem 13
In this section, we show that if $L\in\mathsf{SF}(\mathcal{R})$, then $L\in\mathcal{L}(\mathsf{FC})$.
Proof:
Let us start by proving the following claim:
Claim. For every $L\in\mathsf{SF}(\mathcal{R})$, there exists an $\mathsf{FC}$-formula $\varphi_{L}(x)$ such that $(\mathcal{A}_{w},\sigma)\models\varphi_{L}(x)$ if and only if $\sigma(x)\in L$.
Proof of Claim. Let $L\in\mathsf{SF}(\mathcal{R})$. We proceed by a structural induction (which follows very similarly to the proof of Lemma 5.5 of [1] given in the version with proofs [27]).
  • If $L=\{\mathtt{a}\}$ for some $\mathtt{a}\in\Sigma$, then let $\varphi_{L}(x):=(x\mathbin{\dot{=}}\mathtt{a})$,
  • If $L=w^{*}$ for some $w\in\Sigma^{*}$, then let [^5] We use $z^{p}$ as shorthand for $z$ concatenated with itself $p$-times; which can clearly be expressed in $\mathsf{FC}$.
    $\varphi_{L}(x):=\exists y,z\colon\bigl{(}(x\mathbin{\dot{=}}y\cdot\varrho(w))\land(x\mathbin{\dot{=}}\varrho(w)\cdot y)\land(x\mathbin{\dot{=}}z^{p})\bigr{)},$
    where $p\in\mathbb{N}$ such that $w=\varrho(w)^{p}$. The correctness of $\varphi_{L}(x)$ for this case follows from a result on commuting words: If $uv=vu$ with $u,v\in\Sigma^{*}$, then there exists $n,m\in\mathbb{N}$ and $w\in\Sigma^{*}$ with $u=w^{n}$ and $v=w^{m}$ (See Proposition 1.3.2 of Lothaire [26], and see the proof of Lemma 5.5 of [1] given in [27]).
  • If $L=L_{1}\mathrel{\cup}L_{2}$, then we can assume the existence of $\varphi_{L_{1}}(x)$ and $\varphi_{L_{2}}(x)$ by the induction hypothesis. Then, let $\varphi_{L}(x):=\varphi_{L_{1}}(x)\lor\varphi_{L_{2}}(x)$.
  • If $L=\Sigma^{*}\setminus L_{1}$, then by the induction hypothesis, we assume the existence of $\varphi_{L_{1}}$ and let $\varphi_{L}(x):=\neg\varphi_{L_{1}}(x)$.
  • If $L=L_{1}\cdot L_{2}$, then we can assume the existence of $\varphi_{L_{1}}(x_{1})$ and $\varphi_{L_{2}}(x_{2})$ by the induction hypothesis, with $x$, $x_{1}$ and $x_{2}$ being pairwise distinct. Then, let $\varphi_{L}(x):=\exists x_{1},x_{2}\colon\bigl{(}(x\mathbin{\dot{=}}x_{1}\cdot x_{2})\land\varphi_{L_{1}}(x_{1})\land\varphi_{L_{2}}(x_{2})\bigr{)}$.
This concludes the proof of this claim. $\blacksquare$ (Claim)
Thus, for any $L\in\mathsf{SF}(\mathcal{R})$, we have $\varphi_{L}(x)$ such that $(\mathcal{A}_{w},\sigma)\models\varphi(x)$ if and only if $\sigma(x)\in L$. Now let
$\phi_{L}:=\exists x\colon\Bigl{(}\varphi_{L}(x)\land\forall y,z\colon\bigl{(}((y\mathbin{\dot{=}}x\cdot z)\lor(y\mathbin{\dot{=}}z\cdot x))\rightarrow(z\mathbin{\dot{=}}\varepsilon)\bigr{)}\Bigr{)}.$
The formula $\phi_{L}$ states that there exists a factor $x$ such that $x\in L$, and for any factors $y,z$ with $y=xz$ or $y=zx$, it necessarily holds that $z=\varepsilon$. Consequently, $x$ represents the whole input word. Thus, $\mathcal{A}_{w}\models\phi_{L}$ if and only if $w\in L$.
Appendix B Proof of Lemma 17
In this section, our goal is to establish $\mathsf{SF}(\mathcal{R})=\mathsf{SF}(\mathcal{B}q)$. We start by giving some basic lemmas on bounded regular languages. Recall that $L\subseteq\Sigma^{*}$ is a bounded regular language if $L$ is regular and $L\subseteq w_{1}^{*}\cdots w_{n}^{*}$ for $w_{1},\dots,w_{n}\in\Sigma^{+}$ and $n\geqslant 1$.
Lemma 38.
Any bounded regular language $L\subseteq\Sigma^{*}$ belongs to $\mathsf{SF}(\mathcal{R})$.
Proof:
From Theorem 1.1 of [25], the class of bounded regular languages is the smallest class that contains; the finite languages, the languages $w^{*}$ for any $w\in\Sigma^{*}$, and is closed under concatenation and finite union. Immediately from the definition of $\mathsf{SF}(\mathcal{R})$, we can see every bounded regular language $L\subseteq\Sigma^{*}$ is in $\mathsf{SF}(\mathcal{R})$.
Lemma 39.
If $L\subseteq w^{*}$, for some $w\in\Sigma^{*}$, is regular and $\mathtt{a}\in\Sigma$, then $\mathtt{a}^{-1}L$ and $L\mathtt{a}^{-1}$ are both bounded regular languages.
Proof:
Let $w\in\Sigma^{*}$. Let $I\subseteq\mathbb{N}$ such that $L=\{w^{i}\mid i\in I\}$ is regular. We shall prove that $\mathtt{a}^{-1}L$ is a bounded regular language ($L\mathtt{a}^{-1}$ follows symmetrically). If $w=\mathtt{b}u$ for some $\mathtt{b}\in\Sigma\setminus\{\mathtt{a}\}$ and $u\in\Sigma^{*}$, then $\mathtt{a}^{-1}L=\emptyset$ and therefore $\mathtt{a}^{-1}L$ is a bounded regular language. If $w=\mathtt{a}u$ for some $u\in\Sigma^{*}$, then $\mathtt{a}^{-1}L=u\cdot\{w^{i-1}\mid i\in I\text{ and }i>0\}$. Thus, $\mathtt{a}^{-1}L\subset u^{*}w^{*}$ and thus is a bounded language. Furthermore, if $L^{\prime}$ is regular and $v\in\Sigma^{*}$, then $v^{-1}L$ is also regular (see Theorem 4.5 of [36]). Consequently, $\{w^{i-1}\mid i\in I\text{ and }i>0\}$ is regular, and thus $\mathtt{a}^{-1}L$ is bounded and regular.
B-A Actual proof of Lemma 17
Proof:
By the definition of $\mathcal{R}$ and $\mathcal{B}q$, we have that $\mathcal{R}\subseteq\mathcal{B}q$ and thus we immediately get $\mathsf{SF}(\mathcal{R})\subseteq\mathsf{SF}(\mathcal{B}q)$. To see that $\mathsf{SF}(\mathcal{B}q)\subseteq\mathsf{SF}(\mathcal{R})$, we show that every language $L\in\mathcal{B}q$ belongs to $\mathsf{SF}(\mathcal{R})$. As stated earlier, $\emptyset,\Sigma^{*}\in\mathsf{SF}(\mathcal{R})$, see Section III-A. The regular language $L\subseteq w^{*}$, for any $w\in\Sigma^{*}$, is a bounded regular language, thus $L\in\mathsf{SF}(\mathcal{R})$ (recall Lemma 38). Furthermore, $\mathsf{SF}(\mathcal{R})$ is closed under union and complement (by definition). It therefore remains to show that if $L\in\mathcal{B}q$, then $u^{-1}L,u^{-1}L\in\mathsf{SF}(\mathcal{R})$ for any $u\in\Sigma^{*}$.
The rest of this proof follows closely to the proof of Proposition 3.2 in [14]. We proceed by induction along the length of $u$. If $|u|=0$, then $u^{-1}L=Lu^{-1}=L$ and therefore we are done. For the inductive step, we only consider $u^{-1}L$ as $Lu^{-1}$ follows symmetrically. Considering when $|u|>0$, we have $u^{-1}L=w^{-1}(\mathtt{a}^{-1}L)$ where $u=w\mathtt{a}$ and $\mathtt{a}\in\Sigma$. Thus, by the induction hypothesis, $u^{-1}L\in\mathsf{SF}(\mathcal{R})$ if $\mathtt{a}^{-1}L\in\mathsf{SF}(\mathcal{R})$. To show $\mathtt{a}^{-1}L\in\mathsf{SF}(\mathcal{R})$, we use a structural sub-induction. Clearly $\mathtt{a}^{-1}\emptyset=\emptyset$ and $\mathtt{a}^{-1}\Sigma^{*}=\Sigma^{*}$ and therefore, these cases are trivial. The language $\mathtt{a}^{-1}L$ where $L\subseteq v^{*}$, where $L$ is regular and $v^{*}\in\Sigma^{*}$, is a bounded regular language (see Lemma 39) and therefore $\mathtt{a}^{-1}L\in\mathsf{SF}(\mathcal{R})$, see Lemma 38. Consider $\mathtt{a}^{-1}L$ where $L=L_{1}\mathrel{\cup}L_{2}$. By the sub-induction hypothesis, $\mathtt{a}^{-1}L_{1}\in\mathsf{SF}(\mathcal{R})$ and $\mathtt{a}^{-1}L_{2}\in\mathsf{SF}(\mathcal{R})$ and $\mathtt{a}^{-1}L=\mathtt{a}^{-1}L_{1}\mathrel{\cup}\mathtt{a}^{-1}L_{2}$ and thus we are done with union. For complement: Let $L=\Sigma^{*}\setminus K$ where $K\in\mathcal{B}q$. Note that
$\mathtt{a}^{-1}(\Sigma^{*}\setminus K)=\{v\mid\mathtt{a}v\in\Sigma^{*}\setminus K\}=\\ \{v\mid\mathtt{a}v\in\Sigma^{*}\text{ and }\mathtt{a}v\notin K\}=\Sigma^{*}\setminus\mathtt{a}^{-1}K.$
By the sub-induction hypothesis $\mathtt{a}^{-1}K\in\mathsf{SF}(\mathcal{R})$, and thus we are done with complement. This completes the proof that $\mathcal{B}q\subseteq\mathsf{SF}(\mathcal{R})$, which implies $\mathsf{SF}(\mathcal{B}q)\subseteq\mathsf{SF}(\mathcal{R})$.
Appendix C Proof of Lemma 33
Let us recall Lemma 33:
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a bifix morphism where $h(\mathtt{0})$ is primitive and $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$. Let $\bar{w}_{A}\in\{\mathtt{0},\mathtt{1}\}^{*}$ be the canonical encoding of a set $A\subset\mathbb{N}$ where $\mathsf{min}(A)\geqslant 2|h(\mathtt{1})|$. Then, for any $u\sqsubseteq h(\bar{w}_{A})$ where $|u|>10\cdot\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)$, the pre-image core of $u$ is unique.
Proof:
Let $\mathsf{m}:=\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)$. Working towards a contradiction, assume that $y,y^{\prime}\sqsubseteq\bar{w}_{A}$ are both pre-image cores of $u\sqsubseteq h(\bar{w}_{A})$ where $y\neq y^{\prime}$. Then, $(x,y,z)$ and $(x^{\prime},y^{\prime},z^{\prime})$ are core factorizations of $u$ with respect to $\bar{w}_{A}$ and $h$. It follows that $x\cdot h(y)\cdot z=x^{\prime}\cdot h(y^{\prime})\cdot z^{\prime}$. Let $y=y_{1}\cdot y_{2}\cdots y_{n}$ where $y_{i}\in\{\mathtt{0},\mathtt{1}\}^{*}$ for $i\in[n]$. Likewise, let $y^{\prime}=y_{1}^{\prime}\cdot y_{2}^{\prime}\cdots y_{m}^{\prime}$ where $y^{\prime}_{i}\in\{\mathtt{0},\mathtt{1}\}^{*}$ for $i\in[m]$.
Case 1, $|x|=|x^{\prime}|$
We immediately know that $x=x^{\prime}$ which implies $h(y)\cdot z=h(y^{\prime})\cdot z^{\prime}$. Note that since $h$ is injective, if $z=z^{\prime}$ we have that $y=y^{\prime}$ must hold. Therefore, $z\neq z^{\prime}$ must hold. Without loss of generality, assume that $|z|<|z^{\prime}|$. Thus, $z^{\prime}=z_{2}\cdot z$ for some $z_{2}\in\Sigma^{+}$. That gives us $h(y)\cdot z=h(y^{\prime})\cdot z_{2}\cdot z$ and hence $h(y)=h(y^{\prime})\cdot z_{2}.$ Note that $z_{2}\sqsubset_{\mathsf{pref}}z^{\prime}\sqsubset_{\mathsf{pref}}h(c)$ for some $c\in\{\mathtt{0},\mathtt{1}\}$. Now, if $y^{\prime}\sqsubset_{\mathsf{pref}}y$, then $y=y^{\prime}\cdot y^{\prime\prime}$ for some $y^{\prime\prime}\in\{\mathtt{0},\mathtt{1}\}^{+}$. Hence, $h(y^{\prime})\cdot h(y^{\prime\prime})=h(y^{\prime})\cdot z_{2}$ which implies $h(y^{\prime\prime})=z_{2}$ where $z_{2}\sqsubset_{\mathsf{pref}}h(c)$ for some $c\in\{\mathtt{0},\mathtt{1}\}$. Let $c^{\prime}\in\{\mathtt{0},\mathtt{1}\}$ be the first terminal symbol of $y^{\prime\prime}$. It follows that $c^{\prime}\neq c$ since $z_{2}\sqsubset_{\mathsf{pref}}h(c)$. However, then $h(c^{\prime})\sqsubset_{\mathsf{pref}}h(c)$ as $h(c^{\prime})\sqsubseteq_{\mathsf{pref}}h(y^{\prime\prime})$. This contradicts $h$ being a bifix morphism. Thus, since $y^{\prime}\sqsubset_{\mathsf{pref}}y$ cannot hold, let $\lambda$ be the longest shared prefix of $y$ and $y^{\prime}$. We get that for some $\lambda_{1},\lambda_{2}\in\{\mathtt{0},\mathtt{1}\}^{*}$, the following equality holds:
$h(\lambda)\cdot h(c)\cdot h(\lambda_{1})=h(\lambda)\cdot h(c^{\prime})\cdot h(\lambda_{2})\cdot z_{2}$
where $c,c^{\prime}\in\{\mathtt{0},\mathtt{1}\}$ and $c\neq c^{\prime}$. Thus
$h(c)\cdot h(\lambda_{1})=h(c^{\prime})\cdot h(\lambda_{2})\cdot z_{2}$
However, $h(c)$ is not a prefix of $h(c^{\prime})$ and vice versa; and $h(c)\neq h(c^{\prime})$. Therefore, we have reached a contradiction as the above equality cannot hold.
Case 2, $|x|\neq|x^{\prime}|$
Without loss of generality, assume that $|x|<|x^{\prime}|$. Thus, we have that $x^{\prime}=x\cdot x_{2}$ for some $x_{2}\in\Sigma^{+}$. Therefore, $x\cdot h(y)\cdot z=x\cdot x_{2}\cdot h(y^{\prime})\cdot z^{\prime}$ which implies $h(y)\cdot z=x_{2}\cdot h(y^{\prime})\cdot z^{\prime}$.
Now let $i_{1},i_{2},\dots,i_{n}\in\mathbb{N}$ such that $|h(y_{1}\cdots y_{r})|=i_{r}$ for $r\in[n]$, and let $j_{1},j_{2},\dots,j_{m}\in\mathbb{N}$ such that $|x_{2}\cdot h(y^{\prime}_{1}\cdots y^{\prime}_{p})|=j_{p}$ for $p\in[m]$.
We now argue that $\{i_{1},\dots,i_{n}\}\mathrel{\cap}\{j_{1},\dots,j_{m}\}=\emptyset$. Working towards a contradiction, assume that $i_{r}\in\{j_{1},\dots,j_{m}\}$ for some $r\in[n]$ and further assume $i_{r}$ is the smallest such value. Thus, there exists $j_{p}\in\{j_{1},\dots,j_{m}\}$ such that $i_{r}=j_{p}$. We can therefore write $h(y_{1}\cdots y_{r})=x_{2}\cdot h(y^{\prime}_{1}\cdots y^{\prime}_{p})$. However, since we assumed $i_{r}$ is the smallest such values, it follows that $h(y_{r})\neq h(y^{\prime}_{p})$ and hence either $h(y_{r})$ is a suffix of $h(y^{\prime}_{p})$ or vice versa – depending on whether $|h(y_{r})|<|h(y^{\prime}_{p})|$ or $|h(y^{\prime}_{p})|<|h(y_{r})|$ holds. This contradicts $h$ being bifix and therefore we continue the proof with the knowledge that $\{i_{1},\dots,i_{n}\}\mathrel{\cap}\{j_{1},\dots,j_{m}\}=\emptyset$.
For any $i_{r}\in\{i_{1},\dots,i_{n-1}\}$ such that $y_{r}=y_{r+1}=\mathtt{0}$, we say that $i_{r}$ is $c$-covered for some $c\in\{\mathtt{0},\mathtt{1}\}$ if there is some $p\in[m-1]$ such that $j_{p}<i_{r}<j_{p+1}$ and $y^{\prime}_{p}=c$. If we also need to refer to $j_{p}$, then we say $i_{r}$ is $c$-covered by $j_{p}$. See Fig. 3 for an illustration. We define those points $j_{s}\in\{j_{1},\dots,j_{m-1}\}$ where $y^{\prime}_{s}=y^{\prime}_{s+1}=\mathtt{0}$ as being $c$-covered symmetrically.
Note that if there is some $i_{r}$ such that $y_{r}=y_{r+1}=\mathtt{0}$ which is $\mathtt{0}$-covered, then $h(\mathtt{0})$ is an internal factor of $h(\mathtt{0})\cdot h(\mathtt{0})$. This contradicts our assumption that $h(\mathtt{0})$ is primitive (recall Lemma 28). Likewise, if there is some $j_{s}$ such that $y^{\prime}_{s}=y^{\prime}_{s+1}=\mathtt{0}$ which is $\mathtt{0}$-covered, then $h(\mathtt{0})$ is imprimitive.
[…]
Figure 3: Illustration of $i_{r}$ being $\mathtt{1}$-covered by $j_{p}$.
Recall that $h(y_{1}\cdot y_{2}\cdots y_{n})\cdot z=x_{2}\cdot h(y_{1}^{\prime}\cdot y^{\prime}_{2}\cdots y^{\prime}_{m})\cdot z^{\prime}$. Since $\mathsf{min}(A)>2|h(1)|$, it follows that if $y_{i}=\mathtt{1}$, then $y_{j}=\mathtt{0}$ for $i+1\leqslant j\leqslant\mathsf{min}(i+2|h(1)|+2,n)$; the analogous holds for $y^{\prime}$. For intuition, to arrive at a contradiction, we shall use the fact in the previous sentence to show that there must exist some element of $j_{r}\in\{j_{1},\dots,j_{m-1}\}$ such that $y_{r}^{\prime}=y_{r+1}^{\prime}=\mathtt{0}$ and $j_{r}$ is $\mathtt{0}$-covered – which contradicts $h(\mathtt{0})$ being primitive:
Let $j_{s}\in\{j_{1},\dots,j_{m}\}$ be the smallest value such that $y_{s}^{\prime}=y_{s+1}^{\prime}=\mathtt{0}$. Note that due to the definition of $\bar{w}_{A}$, we have that $s\leqslant 3$ (since if $y_{1}^{\prime}=\mathtt{1}$ then $y_{2}^{\prime}=y_{3}^{\prime}=\mathtt{0}$, and if $y_{2}^{\prime}=\mathtt{1}$ then $y_{3}^{\prime}=y_{4}^{\prime}=\mathtt{0}$). We know that $j_{s}$ must be $\mathtt{1}$-covered (otherwise $h(\mathtt{0})$ is imprimitive). Therefore, there exists some $i_{p}$ such that $i_{p-1}<j_{s}<i_{p}$ and $y_{p}=\mathtt{1}$. Furthermore, due to the fact that $\mathsf{min}(A)>2|h(1)|$, we know $y_{p+1}=y_{p+2}=\dots=y_{p+2|h(1)|+2}=\mathtt{0}$. Now, since $i_{p+1}$ must be $\mathtt{1}$-covered, there exists some $j_{q-1}<i_{p+1}<j_{q}$ such that $y_{q}^{\prime}=\mathtt{1}$. However, $y_{q+1}^{\prime}=y_{q+2}^{\prime}=\mathtt{0}$ and $y_{q+1}^{\prime}$ must be $\mathtt{0}$-covered due to the fact that $i_{p+1}<j_{q}$ and $i_{p+2|h(\mathtt{1})|+2}>j_{q+2}$, along with the fact that $y_{p+1}=y_{p+2}=\dots=y_{p+2|h(\mathtt{1})|+2}=\mathtt{0}$. This contradicts $h(\mathtt{0})$ being primitive and therefore our assumption that there exists two pre-image cores of $u$ must be incorrect.
To conclude the proof of Case 2 and thus the proof of Lemma 33, we show that the size bounds on $|u|$ are enough to achieve the derived contradiction. Recall that $\mathsf{m}:=\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)$. Due to the definition of a core factorization, we know that $|x|<\mathsf{m}$ and $|x^{\prime}|<\mathsf{m}$. Therefore, $j_{s}\leqslant 4\mathsf{m}$ as we established earlier that $s\leqslant 3$. Note that $j_{s}$ is $\mathtt{1}$-covered by $i_{p}$ and therefore $i_{p}\leqslant j_{s}+\mathsf{m}\leqslant 5\mathsf{m}$. Furthermore $i_{p+1}\leqslant 6\mathsf{m}$ and is $1$-covered by $j_{q}$ and therefore $j_{q}\leqslant 7\mathsf{m}$. Finally, $j_{q}$ and $j_{q+1}$ is where the contradiction occurs as $y_{q+1}^{\prime}=y_{q+2}^{\prime}=\mathtt{0}$ and $y_{q+1}^{\prime}$ must be $\mathtt{0}$-covered and $j_{q+1}\leqslant 8\mathsf{m}$. Consequenly, $|u|>9\mathsf{m}$ is enough for a contradiction as $|z|<\mathsf{m}$.
Appendix D Proof of Lemma 35
Let us recall Lemma 35:
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a bifix morphism where $h(\mathtt{0})$ is primitive, and $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$. If $\bar{w}_{A}\equiv_{k+20}\bar{w}_{B}$ for sets $A,B\subseteq\mathbb{N}$ where $\mathsf{min}(A\mathrel{\cup}B)>2\cdot\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)+10$, then $h(\bar{w}_{A})\equiv_{k}h(\bar{w}_{B})$.
Proof:
Let $\Sigma$ be any alphabet where $|\Sigma|\geqslant 2$. Let $h\colon\{\mathtt{0},\mathtt{1}\}^{*}\rightarrow\Sigma^{*}$ be a bifix morphism where $h(\mathtt{0})$ is primitive, and $\varrho(h(\mathtt{1}))\neq h(\mathtt{0})$. Let $\mathsf{m}:=\mathsf{max}(|h(\mathtt{0})|,|h(\mathtt{1})|)$. Assume that $\bar{w}_{A}\equiv_{k+20}\bar{w}_{B}$ for sets $A,B\subseteq\mathbb{N}$ where $\mathsf{min}(A\mathrel{\cup}B)>2\mathsf{m}+10$.
We write $\mathcal{G}$ for the $k$-round game over $h(\bar{w}_{A})$ and $h(\bar{w}_{B})$. Let $\mathcal{G}^{\prime}$ be a $k+20$-round game over $\bar{w}_{A}$ and $\bar{w}_{B}$. Using the fact that Duplicator plays $\mathcal{G}^{\prime}$ using their winning strategy, we shall now show that $h(\bar{w}_{A})\equiv_{k}h(\bar{w}_{B})$.
$u_{i}\cdot h(c_{i})\cdot u_{i}^{\prime}$$c_{i}$$d_{i}$$u_{i}\cdot h(d_{i})\cdot u_{i}^{\prime}$$\mathcal{G}^{\prime}$$\mathcal{G}$
Figure 4: Duplicator’s strategy when Spoiler chooses some $u$ where $|u|>10\mathsf{m}$.
Duplicator’s Strategy
Without loss of generality, let Spoiler chose some $a_{i}\sqsubseteq h(\bar{w}_{A})$ in round $i$. Now we define Duplicator’s response:
  1. 1.
    If $|a_{i}|>10\mathsf{m}$, then let $(u_{i},c_{i},u_{i}^{\prime})$ be the core factorization of $a_{i}$. Let Spoiler choose $c_{i}$ in round $i$ of $\mathcal{G}^{\prime}$ and let $d_{i}$ be Duplicator’s response in $\mathcal{G}^{\prime}$. Then, Duplicator’s response in $\mathcal{G}$ is $b_{i}:=u_{i}\cdot h(d_{i})\cdot u_{i}^{\prime}$. See Fig. 4.
  2. 2.
    If $|a_{i}|\leqslant 10\mathsf{m}$ then let Duplicator respond with $b_{i}:=a_{i}$. We assume that Spoiler chooses $\varepsilon$ in $\mathcal{G}^{\prime}$ for round $i$; the structure Spoiler chooses here is not important.
The case where Spoiler chose some $b_{i}\sqsubseteq h(\bar{w}_{B})$ in round $i$ is defined symmetrically.
Let $\kappa=k+|\Sigma|+1$. Let $\vec{a}=(a_{1},\dots,a_{\kappa})$ and $\vec{b}=(b_{1},\dots,b_{\kappa})$ be the resulting tuples from $\mathcal{G}$ where Duplicator plays their defined strategy. Let $\vec{c}=(c_{1},\dots,c_{\kappa})$ and $\vec{d}=(d_{1},\dots,d_{\kappa})$ be the resulting tuples from the first $k$-round of $\mathcal{G}^{\prime}$ where Duplicator plays their winning $k+20$-round strategy.
Note that Duplicator can still survive $\mathcal{G}^{\prime}$ for an extra $20$ rounds, and therefore must play accordingly. We shall utilize this fact that ensure Duplicator’s strategy for $\mathcal{G}$ is indeed a winning strategy. That is, we will look at possible choices for Spoiler in $\mathcal{G}^{\prime}$ for rounds $k+1,\dots,k+20$ and show that since Duplicator must win $\mathcal{G}^{\prime}$, this enforces Duplicator plays a way which translates to winning $\mathcal{G}$. For a round $k+i$ of $\mathcal{G}^{\prime}$ where $i\geqslant 1$, we denote the choice from $\bar{w}_{A}$ as $c_{\kappa+i}$ and the choice from $\bar{w}_{B}$ as $d_{\kappa+i}$. As an example, from Lemma 34, we know that if $|c_{i}|\leqslant 20$ for any $i\in[k]$, then $c_{i}=d_{i}$ must hold. Likewise, if $|d_{i}|\leqslant 20$ for any $i\in[k]$, then $d_{i}=c_{i}$ must hold.
Correctness
In this section, we shall first prove that Duplicator’s strategy is well-defined. Then, we shall show that it is indeed a winning strategy.
Clearly, for any choice Spoiler makes, there is a unique response defined for Duplicator, see Lemma 33. Thus, to show that Duplicator’s strategy is well-defined, we must now show that Duplicator always responds with a factor of the corresponding word. Assume without loss of generality, that in round $i$ of $\mathcal{G}$ Spoiler chooses $a_{i}\sqsubseteq h(\bar{w}_{A})$. We shall now prove that $b_{i}\sqsubseteq h(\bar{w}_{B})$ does indeed hold, where $b_{i}$ is Duplicator’s response as per their defined strategy.
First we show that if $a_{i}\leqslant 10\mathsf{m}$ then $a_{i}\sqsubseteq h(\bar{w}_{B})$. Let $u\sqsubseteq\bar{w}_{A}$ where $|u|\leqslant 10$, then either:
  • $u=\mathtt{0}^{r}$ where $r\leqslant 10$ or
  • $u=\mathtt{0}^{r}\cdot\mathtt{1}\cdot\mathtt{0}^{s}$, where $r+s<10$.
Since $\mathsf{min}(A)>2\mathsf{m}+10$ and $\mathsf{min}(B)>2\mathsf{m}+10$ we have that $u\sqsubseteq\bar{w}_{B}$. Thus, for such an $a_{i}$, we can find some $u\sqsubseteq\bar{w}_{A}$ such that $a_{i}\sqsubseteq h(u)$ which implies $a_{i}\sqsubseteq h(\bar{w}_{B})$.
Next consider $a_{i}>10\mathsf{m}$. Then $(w_{i},c_{i},w_{i}^{\prime})$ is the core factorization of $a_{i}$, where $c_{i}$ is the $i$-th element chosen from $\bar{w}_{A}$ in $\mathcal{G}^{\prime}$. We have that $(w_{i},d_{i},w_{i}^{\prime})$ is the core factorization of $b_{i}$, where $d_{i}$ is the $i$-th element chosen from $\bar{w}_{B}$ in $\mathcal{G}^{\prime}$. Note that by the definition of the core factorization, there exists $\bar{w}_{i}\sqsubset_{\mathsf{pref}}h(c)$ and $\bar{w}_{i}^{\prime}\sqsubset_{\mathsf{suff}}h(c^{\prime})$ for some $c,c^{\prime}\in\{\mathtt{0},\mathtt{1}\}$ such that
$\bar{w}_{i}\cdot w_{i}\cdot h(c_{i})\cdot w_{i}^{\prime}\cdot\bar{w}_{i}^{\prime}=h(c)\cdot h(c_{i})\cdot h(c^{\prime}),$
where $\bar{w}_{i}\cdot w_{i}=h(c)$ and $w_{i}^{\prime}\cdot\bar{w}_{i}^{\prime}=h(c^{\prime})$.
Now, we look at various rounds of $\mathcal{G}^{\prime}$ after round $k$.
  • In round $k+1$ of $\mathcal{G}^{\prime}$, Spoiler can choose $c_{\kappa+1}=c\cdot c_{i}$, Duplicator must respond with $d_{\kappa+1}=c\cdot d_{i}$.
  • In round $k+2$ of $\mathcal{G}^{\prime}$, Spoiler can choose $c_{\kappa+2}=c_{\kappa+1}\cdot c^{\prime}$, Duplicator must respond with $d_{\kappa+2}=d_{\kappa+1}\cdot c^{\prime}$.
Thus, $d_{\kappa+2}\sqsubseteq\bar{w}_{B}$ where $d_{\kappa+2}=c\cdot d_{i}\cdot c^{\prime}$. Hence, $b_{i}\sqsubseteq h(z_{1}\cdot d_{i}\cdot z_{2})\sqsubseteq h(\bar{w}_{B})$ which implies $b_{i}\sqsubseteq h(\bar{w}_{B})$.
We have proven that Duplicator’s strategy is well defined. Our next focus is to prove that Duplicator’s strategy is a winning strategy. The following claim is sufficient for showing Duplicator’s strategy is a winning strategy.
Claim. For all $l,i,j\in[\kappa]$ we have $a_{l}=a_{i}\cdot a_{j}$ if and only if $b_{l}=b_{i}\cdot b_{j}$.
Proof of Claim. Assume, without loss of generality, that $a_{l}=a_{i}\cdot a_{j}$ for some $l,i,j\in[\kappa]$. We shall show that $b_{l}=b_{i}\cdot b_{j}$. The case where $b_{l}=b_{i}\cdot b_{j}$ implies $a_{l}=a_{i}\cdot a_{j}$ for $l,i,j\in[\kappa]$ follows symmetrically.
To prove $a_{l}=a_{i}\cdot a_{j}$ implies $b_{l}=b_{i}\cdot b_{j}$, we use the following case distinction.
  • Case 1: $|a_{i}|>10\mathsf{m}$ and $|a_{j}|>10\mathsf{m}$.
  • Case 2: $|a_{i}|>10\mathsf{m}$ and $|a_{j}|\leqslant 10\mathsf{m}$.
  • Case 3: $|a_{i}|\leqslant 10\mathsf{m}$ and $|a_{j}|>10\mathsf{m}$.
  • Case 4: $|a_{i}|\leqslant 10\mathsf{m}$ and $|a_{j}|\leqslant 10\mathsf{m}$.
Case 1, $|a_{i}|>10\mathsf{m}$ and $|a_{j}|>10\mathsf{m}$: For $\gamma\in\{i,j,l\}$, let $a_{\gamma}=w_{\gamma}\cdot h(c_{\gamma})\cdot w_{\gamma}^{\prime}$ where $c_{\gamma}\sqsubseteq\bar{w}_{A}$ is the pre-image core of $a_{\gamma}$. As $a_{l}=a_{i}\cdot a_{j}$, we have
$\underbrace{w_{l}\cdot h(c_{l})\cdot w_{l}^{\prime}}_{a_{l}}=\underbrace{w_{i}\cdot h(c_{i})\cdot w_{i}^{\prime}}_{a_{i}}\cdot\underbrace{w_{j}\cdot h(c_{j})\cdot w_{j}^{\prime}}_{a_{j}}.$
We can take $w_{l}\cdot h(c_{l})\cdot w_{l}^{\prime}$ and factorize it as
$a_{l}=w_{l}\cdot\underbrace{h(c_{l,p})\cdot w_{l,p}\cdot w_{l,s}\cdot h(c_{l,s})}_{h(c_{l})}\cdot w_{l}^{\prime},$ (2)
where $w_{l,p}\cdot w_{l,s}=h(\lambda_{l})$ for some $\lambda_{l}\in\{\mathtt{0},\mathtt{1}\}$, and where $c_{l}=c_{l,p}\cdot\lambda_{l}\cdot c_{l,s}$. Furthermore, with the appropriate factorization, it follows that $(w_{l},h(c_{l,p}),w_{l,p})$ is a core factorization of $a_{i}$. That is:
$a_{l}=\underbrace{w_{l}\cdot h(c_{l,p})\cdot w_{l,p}}_{a_{i}}\cdot\underbrace{w_{l,s}\cdot h(c_{l,s})\cdot w_{l}^{\prime}.}_{a_{j}}$
Since $|a_{i}|>10\mathsf{m}$, there is a unique core factorization, see Lemma 33, and thus $w_{l}=w_{i}$, $c_{l,p}=c_{i}$, and $w_{l,p}=w_{i}^{\prime}$. Using the symmetric reasoning, we know $w_{l,s}=w_{j}$, $c_{l,s}=c_{j}$, and $w_{l}^{\prime}=w_{j}^{\prime}$.
Now consider $b_{i}\cdot b_{j}$. By the definiton of $\mathcal{G}^{\prime}$ and Duplicator’s strategy, we know that
$b_{i}\cdot b_{j}=w_{i}\cdot h(d_{i})\cdot w_{i}^{\prime}\cdot w_{j}\cdot h(d_{j})\cdot w_{j}^{\prime}.$
We also know that $w_{i}^{\prime}\cdot w_{j}=h(\lambda_{l})$, and therefore
$b_{i}\cdot b_{j}=w_{i}\cdot h(d_{i})\cdot h(\lambda_{l})\cdot h(d_{j})\cdot w_{j}^{\prime}.$
Likewise, we know that $w_{i}=w_{l}$ and $w_{j}^{\prime}=w_{l}^{\prime}$. Hence
$b_{i}\cdot b_{j}=w_{l}\cdot h(d_{i})\cdot h(\lambda_{l})\cdot h(d_{j})\cdot w_{l}^{\prime}.$
As per the definition of the $\mathcal{G}^{\prime}$ and Duplicator’s strategy, we know that
$b_{l}=w_{l}\cdot h(d_{l})\cdot w_{l}^{\prime}.$
Thus, if $d_{l}=d_{i}\cdot\lambda_{l}\cdot d_{j}$, then $b_{l}=b_{i}\cdot b_{j}$. Note that $|h(\lambda_{l})|\leqslant\mathsf{m}$, because $\lambda_{l}\in\{\mathtt{0},\mathtt{1}\}$.
In round $k+1$, Spoiler can choose $h(\bar{w}_{A})$ and $c_{\kappa+1}:=c_{i}\cdot\lambda_{l}$. Hence, Duplicator must respond with $d_{\kappa+1}:=d_{i}\cdot\lambda_{l}$, since $\lambda_{l}$ is a constant. Then, in round $k+2$, Spoiler can choose $c_{\kappa+2}:=c_{\kappa+1}\cdot c_{j}$ and Duplicator must respond with $d_{\kappa+2}:=d_{\kappa+1}\cdot d_{j}$. As $c_{\kappa+2}=c_{i}\cdot\lambda_{l}\cdot c_{j}=c_{l}$, it must hold that $d_{\kappa+2}=d_{i}\cdot\lambda_{l}\cdot d_{j}=d_{l}$. Thus, $d_{l}=d_{i}\cdot\lambda_{l}\cdot d_{j}$ and consequently $b_{l}=b_{i}\cdot b_{j}$.
Case 2, $|a_{i}|>10\mathsf{m}$ and $|a_{j}|\leqslant 10\mathsf{m}$: Note that (from the definition of Duplicator strategy) since $|a_{j}|\leqslant 10\mathsf{m}$, we have that $b_{j}=a_{j}$. Since $a_{l}=a_{i}\cdot a_{j}$, we can write
$w_{l}\cdot h(c_{l})\cdot w_{l}^{\prime}=w_{i}\cdot h(c_{i})\cdot w_{i}^{\prime}\cdot a_{j},$
where $(w_{l},c_{l},w_{l}^{\prime})$ is the unique core factorization of $a_{l}$ and $(w_{i},c_{i},w_{i}^{\prime})$ is the unique core factorization of $a_{i}$.
Using the same reasoning as in Case 1 along with Lemma 33, we have that $w_{l}=w_{i}$ and $c_{i}\sqsubseteq_{\mathsf{pref}}c_{l}$. Refer back to Eq. 2 for some intuition. Thus, there is some $\lambda_{l}\sqsubseteq\bar{w}_{A}$ such that
$w_{l}\cdot\underbrace{h(c_{i})\cdot h(\lambda_{l})}_{h(c_{l})}\cdot w_{l}^{\prime}=w_{l}\cdot h(c_{i})\cdot w_{i}^{\prime}\cdot a_{j}.$
Therefore, $h(\lambda_{l})\cdot w_{l}^{\prime}=w_{i}^{\prime}\cdot a_{j}$. Since $w_{i}^{\prime}\leqslant\mathsf{m}$ and $a_{j}\leqslant 10\mathsf{m}$, we know $|h(\lambda_{l})|\leqslant 11\mathsf{m}$, and hence $\lambda_{l}\leqslant 11$. Spoiler can choose $c_{\kappa+1}:=\lambda_{l}$ in round $k+1$ of $\mathcal{G}^{\prime}$, and therefore Duplicator must respond with $d_{\kappa+1}:=\lambda_{l}$, see Lemma 34. Furthermore, in round $k+2$ of $\mathcal{G}^{\prime}$, Spoiler can choose $c_{l}=c_{i}\cdot\lambda_{l}$ and thus $d_{l}=d_{i}\cdot\lambda_{l}$ must hold.
Now let us consider
$b_{l}=w_{l}\cdot h(d_{l})\cdot w_{l}^{\prime}.$
As $d_{l}=d_{i}\cdot\lambda_{l}$, we have
$b_{l}=w_{l}\cdot h(d_{i})\cdot h(\lambda_{l})\cdot w_{l}^{\prime}.$
Due to the fact that $h(\lambda_{l})\cdot w_{l}^{\prime}=w_{i}^{\prime}\cdot a_{j}$ and $b_{j}=a_{j}$, we know $h(\lambda_{l})\cdot w_{l}^{\prime}=w_{i}^{\prime}\cdot b_{j}$. Using this along with the fact that $w_{l}=w_{i}$, we get
$b_{l}=w_{i}\cdot h(d_{i})\cdot w_{i}^{\prime}\cdot b_{j}.$
Due to the definition of the $\mathcal{G}^{\prime}$ and Duplicator’s strategy, we know that $b_{i}=w_{i}\cdot h(d_{i})\cdot w_{i}^{\prime}$ and hence we arrive at $b_{l}=b_{i}\cdot b_{j}$.
Case 3, $|a_{i}|\leqslant 10\mathsf{m}$ and $|a_{j}|>10\mathsf{m}$: This follows symmetrically from Case 2.
Case 4,$|a_{i}|\leqslant 10\mathsf{m}$ and $|a_{j}|\leqslant 10\mathsf{m}$: Note that since $|a_{i}|\leqslant 10\mathsf{m}$ and $|a_{j}|\leqslant 10\mathsf{m}$, we have that $a_{i}=b_{i}$ and $a_{j}=b_{j}$, by the definition of Duplicator’s strategy. If $|a_{l}|\leqslant 10\mathsf{m}$ then it is trivial that $a_{l}=a_{i}\cdot a_{j}$ implies $b_{l}=b_{i}\cdot b_{j}$, as $a_{l}=b_{l}$. Therefore, we continue this case under the assumption that $|a_{l}|>10\mathsf{m}$. It follows that $a_{l}=w_{l}\cdot h(c_{l})\cdot w_{l}^{\prime}$ where $(w_{l},c_{l},w_{l}^{\prime})$ is the core factorization, and $c_{l}$ is the element chosen from $\bar{w}_{A}$ in round $l$ of $\mathcal{G}^{\prime}$. Thus, $b_{l}=w_{l}\cdot h(d_{l})\cdot w_{l}^{\prime}$ where $d_{l}$ is the element chosen from $\bar{w}_{B}$ in round $l$ of $\mathcal{G}^{\prime}$. Since $|a_{i}|\leqslant 10\mathsf{m}$ and $|a_{j}|\leqslant 10\mathsf{m}$ it follows that $|a_{l}|\leqslant 20\mathsf{m}$, and thus $c_{l}<20$. Thus, from Lemma 34, we know that $d_{l}=c_{l}$ and hence $a_{l}=b_{l}$. Consequently, $a_{l}=a_{i}\cdot a_{j}$ implies $b_{l}=b_{i}\cdot b_{j}$. $\blacksquare$ (Claim) Thus, we have given a winning strategy for Duplicator for $\mathcal{G}$ using Duplicator’s winning strategy for $\mathcal{G}^{\prime}$.